### v1

# A novel efficient structure-preserving exponential integrator for Hamiltonian systems

Abstract
We propose a linearly implicit structure-preserving numerical method for semilinear Hamiltonian systems with polynomial nonlinearities, combining Kahan’s method and exponential integrator. This approach efficiently balances computational cost, accuracy and the preservation of key geometric properties, including symmetry and near-preservation of energy. By requiring only the solution of a single linear system per time step, the proposed method offers significant computational advantages while comparing with the state-of-the-art symmetric energy-preserving exponential integrators. The stability, efficiency and long-term accuracy of the method are demonstrated through numerical experiments on systems such as the Henon-Heiles system, the Fermi-Pasta-Ulam system and the two-dimensional Zakharov-Kuznestov equation.
keywords:
structure-preserving, energy preservation, linearly implicit , exponential integrator, symmetric
1 Intorduction
This study is concerned with semilinear systems of the form:
$\dot{x}(t)=Ax(t)+f(x(t)),\quad x(t_{0})=x_{0},$ (1)
where the matrix $A$ often represents a spatial discretization of a linear unbounded differential operator, and its eigenvalues have large negative real parts or are purely imaginary with substantial magnitude. The nonlinear term $f$ satisfies the Lipschitz condition. Such systems appear in a wide range of applications, including mechanics [1], quantum physics [2], charged-particle dynamic [3] and discretizations of partial differential equations (PDEs) [4]. For these types of systems, exponential integrators are commonly employed due to their ability to efficiently handle the linear part through matrix exponentials, allowing for larger step sizes and improved accuracy in stiff and oscillatory problems [5].
Semilinear systems (1) often exhibit Hamiltonian structures like:
$\begin{split}\dot{x}(t)&=Q\nabla H(x(t)),\\ x(t_{0})&=x_{0},\quad H(x(t))=\frac{1}{2}x^{\mathsf{T}}Mx+U(x),\end{split}$ (2)
where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-order part [6]. For such systems, the Hamiltonian remains invariant along the exact solution trajectory. Moreover, when $Q$ is a canonical skew-symmetric matrix, the system preserves the symplectic two-form. Maintaining these geometric structures is essential for the long-time numerical integration of such systems, as it ensures accuracy and stability [7]. To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 8], symplecticity-related exponential integrators [9, 10, 11, 12, 13], volume-preserving exponential integrator [14] and energy-preserving exponential integrators [15, 16]. Extensive studies have demonstrated that these specialized methods deliver significantly enhanced long-term performance compared to conventional high-order schemes.
In recent years, energy-preserving exponential integrators have been widely studied due to their effectiveness in accurately capturing the energy dynamics of Hamiltonian systems [15, 17, 18, 19, 20, 21, 22, 23]. A prominent example is the exponential averaged vector field (EAVF) method [15] which combines the discrete gradient approach with exponential integrator. This hybrid strategy has inspired the development of a broad class of exponential energy-preserving schemes. The computation of such fully implicit methods requires solving nonlinear systems at each time step—a process that can be computationally demanding. As an alternative, linearly implicit structure-preserving methods have been explored. These methods offer similar stability as fully explicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient. One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [24, 25, 18], which conserves a modified energy for semilinear Hamiltonian systems; however they are not symmetric. A number of further developments along these lines can be found in [12, 13, 21, 26, 27]. Another category involves multiple-point methods that portion out the nonlinearity [26, 8]. These methods have the advantage of preserving both a modified energy and symmetry, but they are inherently multistep-even for systems with quadratic vector fields-which may bring stability concerns when compared to single-step method [28].
In this work, we propose a linearly implicit structure-preserving method for semilinear Hamiltonian systems with polynomial nonlinearities, by combining Kahan’s method [29] with exponential integrators. Kahan’s method is employed due to several favorable properties: i) it is linearly implicit, requiring only the solution of a single linear system per time step, which improves computational efficiency compared to fully implicit methods; ii) it is symmetric, a key property for ensuring stable long-term integration; and iii) it preserves a modified energy and measure, geometric characteristics that are also satisfied by the exact flow of the original system. The proposed method inherits all advantageous properties of Kahan’s method except for the exact conservation of a modified energy and measure. However, we show that the energy error remains bounded and exhibits oscillatory behavior over time. Moreover, we derive an explicit analytical expression for the energy error, offering insights into the method’s long-term energy behavior. Through numerical experiments, we demonstrate that the proposed method achieves superior computational efficiency compared to both the fully implicit exponential AVF method and the linearly implicit method based on the polarized discrete gradient. Additionally, we provide a rigorous proof that the proposed method achieves second-order convergence.
The remainder of this paper is organized as follows. In Section 2, we introduce the proposed linearly implicit structure-preserving exponential integrator and discuss its key properties, including symmetry, energy preservation and convergence. In Section 3, we present numerical experiments to demonstrate the stability, computational efficiency, and long-term behavior of the proposed method using different types of differential equations. Finally, conclusions and potential directions for future research are discussed in the last section.
2 The new linearly implicit structure-preserving symmetric method
Before proposing the new exponential integrator, we will introduce Kahan’s method. For a Hamiltonian system with a cubic Hamiltonian, Kahan’s method has the form
$x_{n+1}=Q(-\frac{1}{2}\nabla H(x_{n})+2\nabla H(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla H(x_{n+1})),$
and it is shown to preserve a modified energy as well as a measure in [29]. In particular, if the energy function is homogeneous, the modified energy can be written as $\bar{H}=C(x_{n},x_{n},x_{n+1})$, where $C$ is a symmetric trilinear form defined by the Hessian of the energy function $H$
$C(x,y,z)=\frac{1}{6}x^{\mathsf{T}}H^{\prime\prime}(y)z.$
When $H$ is nonhomogeneous, we can always add an extra auxiliary variable to transfer the problem to a homogeneous setting. In addition, Kahan’s method is linearly implicit. These properties guarantee the superior numerical behavior of Kahan’s method during long-time integration [28].
For a semilinear system (1), we consider the following variation of constants formula on interval $[t_{0},t_{0}+h]$
$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$ (3)
The key idea of constructing exponential integrator is to integrate the linear part of equation (3) exactly meanwhile finding an approximation to the integration on the right-hand side. For example, the exponential Euler method is obtained by approximating the nonlinear term by $f(x(t_{0}))$ [5], thereby having the form
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)f(x_{n}),$
where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field method is obtained by approximating the nonlinear term by the discrete gradient [15], possessing the form
[…]
In this work, we approximate the nonlinear term using Kahan’s method, leading to the new linearly implicit exponential integrator suitable for solving semilinear systems (1):
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)(-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})).$ (4)
We denote this newly derived scheme as the EKahan scheme, and its properties are described in detail below.
Lemma 1.
The EKahan scheme (4) is symmetric.
Proof.
[…]
$x_{n}=e^{-Ah}x_{n+1}-h\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (5)
Multiplying both sides of (5) by $e^{hA}$, we get
$e^{hA}x_{n}=x_{n+1}-he^{hA}\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (6)
Using the identiy $\phi(-V)=e^{-V}\phi(V)$ for any matrix $V$, we can rewrite the above equation (6) as
$e^{hA}x_{n}=x_{n+1}-h\phi(hA)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$
This equation is equivalent to (4), confirming that the EKahan scheme (4) is symmetric. ∎
Lemma 2.
The EKahan scheme (4) is second-order.
Proof.
By applying Taylor expansion to $f(\frac{x_{n}+x_{n+1}}{2})$ and $f(x_{n+1})$ at the point $x_{n}$ and using the fact that $f$ is quadratic, we can obtain
$-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})=f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n}).$
Therefore, EKahan scheme (4) can be rewritten as
$x_{n+1}$ $=e^{hA}x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})].$ (7)
Since the matrix exponential satisfies $e^{hA}=I+h\phi(hA)A$, we rewrite equation (7) as
$x_{n+1}$ $=(I+h\phi(hA)A)x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})]$
$=x_{n}+h\phi(hA)g(x_{n})+\frac{h}{2}\phi(hA)f^{\prime}(x_{n})(x_{n+1}-x_{n}),$ (8)
where the function $g(x)$ is the right-hand side of the semilinear equation (1), i.e., $g(x):=Ax+f(x)$. Supposing the matrix $I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n})$ is invertible, solving for $x_{n+1}-x_{n}$ from equation (2) gives
$x_{n+1}-x_{n}=h(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n}).$
By $L^{\prime}H\hat{o}pital^{\prime}s\;rule$, we have $\lim\limits_{z\rightarrow 0}\frac{e^{z}-1}{z}=1$, so $\lim\limits_{h\rightarrow 0}\phi(hA)=I$ and
$\lim\limits_{h\rightarrow 0}(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n})=g(x_{n}).$
Therefore, we have the following truncation for the EKahan scheme (7)
$x_{n+1}$ $=x_{n}+hg(x_{n})+O(h^{2}).$ (9)
By Taylor expansion, the exact solution of the semilinear Hamiltonian system (1) satisfies
$x(t_{n+1})=x(t_{n})+hg(x(t_{n}))+\frac{h^{2}}{2}(A+f^{\prime}(x(t_{n})))g(x(t_{n}))+O(h^{3}).$ (10)
Supposing $x_{n}$ is exactly solved and subtracting equation (10) from equation (9), we get
$x_{n+1}-x(t_{n+1})=O(h^{2}).$
Moreover, EKahan scheme (4) is symmetric from Lemma 1, confirming that this method is at least second-order. ∎
Theorem 1.
[…]
$\|f(x_{2})-f(x_{1})\|\leq L\|x_{2}-x_{1}\|.$
Then, the error bound
$\|x(t_{n})-x_{n}\|\leq e^{k_{2}T}(k_{1}\|x(t_{0})-x_{0}\|+k_{3}h^{2})$
holds in $0\leq nh\leq T$ for constant $k_{1},\;k_{2},\;k_{3}$.
Proof.
The assumptions about the operator $A$ guarantee the boundness of $\|\phi(hA)\|$ and $\|e^{jhA}\|$, $j=1,2,\dots$, and we denote the bound by $M$; see Lemma 2.4 from [5]. Since EKahan scheme (4) is second-order, we get the following truncation when inserting the exact solutions $x(t_{n})$ and $x(t_{n+1})$ to the numerical scheme:
$\begin{split}x(t_{n+1})=&e^{hA}x(t_{n})+h\phi(hA)[-\frac{1}{2}f(x(t_{n}))-\frac{1}{2}f(x(t_{n+1}))\\ &+2f(\frac{x(t_{n})+x(t_{n+1})}{2})]+O(h^{3}).\end{split}$ (11)
Denote the numerical error terms by $\varepsilon_{n}=x(t_{n})-x_{n}$ and
$\delta_{n+1,n}=$ $-\frac{1}{2}f(x(t_{n}))+2f(\frac{x(t_{n})+x(t_{n+1})}{2})-\frac{1}{2}f(x(t_{n+1}))$
$-\big{(}-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})\big{)}.$
Subtracting equation (4) from equation(11), we get the following error recursion equation
$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+O(h^{3}).$
This recursion equation gives
$\varepsilon_{n}=e^{nAh}\varepsilon_{0}+\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}+\sum\limits_{j=0}^{n-1}e^{jAh}O(h^{3}).$ (12)
Since $f$ satisfies $Lipschitz\;condition$, we have
$\|\delta_{n-j,n-j-1}\|\leq\frac{3}{2}L\|\varepsilon_{n-j}\|+\frac{3}{2}L\|\varepsilon_{n-j-1}\|.$
Then we have
$\|\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}\|\leq\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|.$ (13)
Combing (12) and (13), we obtain
$\|\varepsilon_{n}\|\leq\|e^{TA}\|\|\varepsilon_{0}\|+\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+nMO(h^{3}).$ (14)
For small $h$ such that $1-\frac{3}{2}hLM$ is positive, suppose there exists a positive real number $k_{3}$ satisfying $\frac{nMO(h^{3})}{1-\frac{3}{2}hLM}\leq k_{3}h^{2}$. Denote by $k_{1}=\frac{\|e^{TA}\|\|\varepsilon_{0}\|}{1-\frac{3}{2}hLM}$ and $k_{2}=\frac{3LM^{2}}{1-\frac{3}{2}hLM}$. Then we can obtain the following estimation for the numerical error $\varepsilon_{n}$ from the above relation (14):
$\|\varepsilon_{n}\|$ $\leq k_{1}\|\varepsilon_{0}\|+k_{2}h\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+k_{3}h^{2}$
$\leq e^{nk_{2}h}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2})$
$=e^{k_{2}T}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2}).$
The last inequality follows from the Gronwall’s inequality. ∎
Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variational constant formula applied to the semilinear system (2). The proof of this lemma is inspired by Theorem 1 in [15].
Lemma 3.
For the semilinear system (2), any exponential integrator with the form
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1}),$
satisfies the following equation
$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n+1},x_{n})=0,$
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
Proof.
Assuming that the matrix $M$ is not singular and denoting by $V=hA=hQM$ and ^$\hat{\nabla}U=M\breve{\nabla}U$, we get the following equation after replacing $x_{n+1}$ by $e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1})$:
$\begin{split}&\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}{x_{n}}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1})\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+h{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+\frac{1}{2}h^{2}\hat{\nabla}U^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}-I)\hat{\nabla}U+{\hat{\nabla}U}^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}\hat{\nabla}U\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}(e^{V}-I)^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{\breve{\nabla}U}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U\\ &=\frac{1}{2}(x_{n}+\breve{\nabla}U)^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)(x_{n}+\breve{\nabla}U)+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-Me^{V})\breve{\nabla}U\\ &=0.\end{split}$
The last equation follows from the fact that ${e^{V}}^{\mathsf{T}}Me^{V}-M$ and ${e^{V}}^{\mathsf{T}}M-Me^{V}$ are skew symmetric.
If $M$ is singular, we follow the idea in [15], introducing a series of symmetric and nonsingular matrices $M_{\delta}$ which converges to $M$ when $\delta$ approaches zero to complete the proof. ∎
The discretized gradient ^$\hat{\nabla}U(x_{n},x_{n+1})$ can include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method. We denote the discretized gradient obtained via Kahan’s method by $\nabla U_{\text{K}}$, which can be written as
$\nabla U_{\text{K}}=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
Theorem 2.
For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, EKahan scheme (4) nearly preserves the exact energy and has the following error estimation
$H_{n+1}-H_{n}$ $=U(x_{n+1}-x_{n}),$
where $H_{n}=\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n})$ is the discrete energy at $t=t_{n}$.
Proof.
Based on Lemma 3, we have
$\begin{split}H_{n+1}-H_{n}&=\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n+1})-U(x_{n})\\ &=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).\end{split}$ (15)
For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla U_{\text{K}}=\frac{U^{\prime\prime}(x_{n+1})x_{n}}{2}$.
Then, the above energy error (15) can be reformulated as
$H_{n+1}-H_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}(x_{n+1}-x_{n})^{\mathsf{T}}(U^{\prime\prime}(x_{n+1})x_{n+1}+U^{\prime\prime}(x_{n})x_{n}-2U^{\prime\prime}(x_{n+1})x_{n})$
$=\frac{1}{6}(x_{n+1}-x_{n})^{\mathsf{T}}U^{\prime\prime}(x_{n+1}-x_{n})(x_{n+1}-x_{n})$
$=U(x_{n+1}-x_{n}).$
For Hamiltonian systems (2) with high-order polynomial function $U(x)$, EKahan method can be generalized using the polarization technique presented in [30]. Suppose $U(x)$ is a homogeneous polynomial of order $k+2$, EKahan scheme is given by
$x_{n+k}=e^{khQM}x_{n}+kh\phi(khQM)Q\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k}),$ (16)
where $\nabla U_{\text{K}}(\cdot,\cdots,\cdot)$ is a symmetric $k+1$-linear polarized function of $\nabla U(x)$, satisfying
$\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$
In this case, we give the following corollaries to describe EKahan’s properties.
Corollary 1.
For semilinear system (2) with high-order polynomial function $U(x)$, the EKahan scheme (16) is symmetric and second order if the initial starting points $x_{n+1},\cdots,x_{n+k}$ are given using a at least second-order method.
This corollary is a generalization of Lemma 1 and 2 for Hamiltonian systems (2) with cubic polynomial function $U(x)$ to systems with high-order polynomial functions, and the proof follows similarly.
Corollary 2.
For the semilinear system (2), any exponential integrator with the form
$x_{n+k}=e^{khA}x_{n}+kh\phi(khA)Q\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k}),$
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretiztion of the true gradient $\nabla U(x_{n})$, satisfies the following equation
$\frac{1}{2k}x_{n+k}^{\mathsf{T}}Mx_{n+k}-\frac{1}{2k}x_{n}^{\mathsf{T}}Mx_{n}+\frac{1}{k}(x_{n+k}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})=0.$
The proof of this corollary follows similarly as the proof of Lemma 3.
Corollary 3.
For a homogeneous function $U(x)$, EKahan scheme (16) nearly preserves the exact energy and has the following error estimation:
$H_{n+1}-H_{n}=$ $\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})$
$+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})$
$-2\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}),$
where $H_{n}=\frac{1}{2k}\sum\limits_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$ is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\cdot,\dots,\cdot,\cdot)$ is a symmetric $k+2$ linear function defined via the $k+2$ tensor associated with the homogeneous function $U(x)$ and satisfies $\bar{U}(x,x,\dots,x,x)=U(x)$.
The proof of this corollary proceeds analogously to that of Theorem 2. For polynomial nonlinear functions $U(x)$ of order greater than three, the energy error can no longer be expressed in the simple form presented in Theorem 2; nevertheless, Corollary 3 demonstrates that the energy error over a single time step can be effectively controlled by the corresponding change in the numerical solution, thereby ensuring the boundedness of the energy error.
Remark 1.
For a nonhomogeneous polynomial $U(x)$, we can extend it to a homogeneous function by introducing an auxiliary variable $x^{0}$ to the original vector $x$. Denote the extended variable as $\bar{x}=(x^{0},x^{1},\cdots,x^{m})^{\mathsf{T}}$, where $x^{i}$, $i=1,\cdots,m$, are the components of the vector $x$. The nonhomogeneous cubic polynomial $U(x)$ is extended to the homogeneous function $\bar{U}(x^{0},x^{1},\cdots,x^{m})$ such that $\bar{U}(1,x^{1},\cdots,x^{m})=U(x)$. Furthermore, to preserve the structural properties of the original system in the extended formulation, we also augment the skew-symmetric matrix $Q$ to $\bar{Q}$ and the symmetric matrix $M$ to $\bar{M}$ by introducing an extra row and column of zero elements. Notably, $\bar{Q}$ remains skew-symmetric. Furthermore, applying the EKahan method to the modified differential equation
$\dot{\bar{x}}(t)=\bar{Q}\bar{M}\bar{x}(t)+\bar{Q}\nabla\bar{U}(\bar{x}(t)),\quad\bar{x}(t_{0})=[1,x_{0}^{\mathsf{T}}]^{\mathsf{T}},$
produces the same update for $x$ as applying the EKahan method directly to the original reduced system (2). Therefore, all results previously established for systems with homogeneous polynomial nonlinearities extend naturally to those involving nonhomogeneous polynomial terms.
3 Numerical test
In this section, we apply the proposed EKahan scheme to three systems including an ODE and two PDEs with high dimension and high-order nonlinear functions. We compare EKahan scheme with the state-of-the-art symmetric structure-preserving methods including the following three schemes:
  • Kahan: the Kahan method presented in [29];
  • EAVF: the exponential Averaged Vector Field method presented in [15];
  • LIEEP: the linearly implicit energy-preserving exponential scheme illustrated in [8].
We define the global error by
$\mathcal{E}_{G}=\max\limits_{n\geq 0}\|y_{n}-y(t_{n})\|,$
and the energy error at time $t_{n}=t_{0}+nh$ by
$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert,$
where $y(t_{n})$ is the reference exact solution and $H_{n}$ is the discrete energy defined in Theorem 2. The reference exact solution is obtained using the 6-order continuous Runge-Kutta (CRK) method [31] with the form
$\left\{\begin{array}[]{l}y_{n+1/3}=y_{n}+hQ\int_{0}^{1}\left(\frac{37}{27}-\frac{32}{9}\sigma+\frac{20}{9}\sigma^{2}\right)\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\\ y_{n+2/3}=y_{n}+hQ\int_{0}^{1}\left(\frac{26}{27}+\frac{8}{9}\sigma-\frac{20}{9}\sigma^{2}\right)\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\\ y_{n+1}=y_{n}+hQ\int_{0}^{1}\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\end{array}\right.,$
where
$Y_{\sigma}=$ $-\frac{(3\sigma-1)(3\sigma-2)(\sigma-1)}{2}y_{n}+\frac{3\sigma(3\sigma-2)(3\sigma-3)}{2}y_{n+1/3}$
$-\frac{3\sigma(3\sigma-1)(3\sigma-3)}{2}y_{n+2/3}+\frac{\sigma(3\sigma-1)(3\sigma-2)}{2}y_{n+1},$
and the integrals are evaluated exactly by the 5-point GL quadrature. For both LIEEP method and EKahan scheme (16), multiple initial starting points are required. These starting values are obtained using the aforementioned reference exact solution. All numerical experiments were conducted using MATLAB (R2022b) on a 2020 MacBook Air equipped with an Apple M1 processor and 16 GB of unified memory.
Henon-Heiles system
The Henon-Heiles model [32] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
$\left\{\begin{array}[]{ll}{\ddot{q}_{1}=-q_{1}-2Dq_{1}q_{2}},\\ {\ddot{q}_{2}=-q_{2}-Dq_{1}^{2}+Cq_{2}^{2}},\\ \end{array}\right.$
where the independent variable $t\in[0,T]$. We consider $C=1$ and $D=1$. By introducing the momentum variables $\dot{q}_{1}=p_{1},\;\dot{q}_{2}=p_{2}$, the system can be rewritten in the Hamiltonian form
$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$ (17)
where $y=(q_{1},q_{2},p_{1},p_{2})^{\mathsf{T}}$, $Q$ is the canonical skew-symmetric matrix, $M=I_{4}$ denotes the $4\times 4$ identity matrix, and the potential function is defined as $U(y)=q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}$.
Applying the EKahan scheme to equation (17) gives
$y_{n+1}=e^{hQM}y_{n}-h\phi(hQM)Q[\frac{1}{2}\nabla{U}(y_{n})-2\nabla U(\frac{y_{n+1}+y_{n}}{2})+\frac{1}{2}\nabla U(y_{n+1})],$
and we can get the accurate form of the exponential function
$e^{hQM}=\begin{pmatrix}\cos(h)I_{2}&\sin(h)I_{2}\\ -\sin(h)I_{2}&\cos(h)I_{2}\end{pmatrix}.$
Consider the initial data $y_{0}=(0,-0.082,0,0)^{\mathsf{T}}$, $T=100$ and time step size $h_{i}=0.02/2^{i},\;i=0,1,\dots,4$. Figure 1(a) demonstrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded, varying from $10^{-18}$ to $10^{-7}$. Figure 1(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 1: Energy behaviour for the Henon-Heiles equation with $T=100$ and $h_{0}=0.02$. Left: energy error computed by all schemes; right: the error of the true and deduced energy deviation by EKahan method shown in Theorem 2.
In Figure 2, we consider the global errors at $T=100$ and the computational cost using different step sizes. Figure 2(a) confirms that the EKahan method is second-order, and Figure 2(b) shows that Ekahan method is more efficient compared to the fully implicit EAVF method and the other two types of linearly implicit methods. In addition, we observe from Figure 2(a) that the exponential integrators are more accurate than Kahan’s method since the linear part are exactly integrated using exponential integrators.
(a) order plot
(b) efficiency plot
Figure 2: Global error and computational time of different methods for the Henon-Heiles equation with $T=100$ and different time step sizes.
FPU system with a cubic Hamiltonian
The $\alpha$-Fermi-Pasta-Ulam (FPU) system, originating from the study by Fermi, Pasta, Ulam, and Tsingou, has evolved into a paradigmatic model in nonlinear wave theory [33]. We study a continuum generalization of the FPU system
$\frac{\partial^{2}u}{\partial t^{2}}=\beta\frac{\partial^{3}u}{\partial t\partial^{2}x}+\frac{\partial^{2}u}{\partial x^{2}}[1+\epsilon(\frac{\partial u}{\partial x})^{p}]-\gamma\frac{\partial u}{\partial t}-m^{2}u,$
where the independent variables $(x,t)\in[0,L]\times[0,T]$, the parameter $\epsilon>0$, the internal damping coefficient $\beta\geq 0$, and the external damping coefficient $\gamma\geq 0$. By introducing $\frac{\partial u}{\partial t}=v$, we can reformulate the second-order partial differential equation into the following first-order differential equations
$\left\{\begin{array}[]{ll}{\partial_{t}u=v,}\\ {\partial_{t}v=\beta\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}u}{\partial x^{2}}(1+\epsilon(\frac{\partial u}{\partial x})^{p})-\gamma v-m^{2}u}.\\ \end{array}\right.$
Setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
$\frac{\partial y}{\partial t}=\mathcal{Q}\frac{\delta\mathcal{H}}{\delta y},$
where
$\mathcal{Q}=\left[\begin{array}[]{cc}0&1\\ -1&\beta\partial^{2}_{x}-\gamma\\ \end{array}\right],\quad\mathcal{H}=\int_{0}^{L}H(t,u,v,u_{x})\mathrm{d}x,$
and the local energy density $H(t,u,v,u_{x})$ is given by
$H(t,u,v,u_{x})=\frac{1}{2}u_{x}^{2}+\frac{m^{2}}{2}u^{2}+\frac{1}{2}v^{2}+\epsilon\frac{u_{x}^{p+2}}{(p+1)(p+2)}.$
We consider $p=1$ and homogeneous Dirichlet boundary conditions $u(0,t)=u(L,t)=0$. Employing the second-order central difference to discretize the second derivative $\partial_{x}^{2}$ and the forward difference to discretize $\partial_{x}$, we derive the semi-discrete system
$\dot{y}=Q(My+\nabla U(y)),$
where
$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u_{j+1}-u_{j}}{\Delta x})^{3}.$
We set $m=0$, $\epsilon=\frac{3}{4}$, $L=128$, $T=100$, the spatial discretization to be $\Delta x=1$ and the time step size to be $h_{i}=1/2^{i}$, $i=1,2,\ldots,4$. The initial conditions are given by
$u_{j}(0)=q_{j}(0),\quad v_{j}(0)=\dot{q}_{j}(0),$
where the function $q_{j}(t)$ takes the form
$q_{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
with $\alpha=0.1$.
We consider three different settings of damping coefficients: i) a conservative system with $\gamma=0$ and $\beta=0$, ii) a dissipative system with $\gamma=0.1$ and $\beta=0$, and iii) a dissipative system with $\gamma=0$ and $\beta=2$. Figure 3 shows that EKahan method preserves the energy structure of the equation in both conservative and dissipative settings:
  1. i)
    the energy error of all methods stays bounded and oscillated in the undamped case in Figure 3(a),
  2. ii)
    the per-step energy deviation under EKahan discretization shown in Theorem 2 is validated in Figure 3(b) ,
  3. iii)
    the energy of all methods drop along time for systems with a positive internal or external damping coefficient in Figure 3(c) and 3(d).
(a) $\gamma=0$, $\beta=0$
(b) $\gamma=0$, $\beta=0$
(c) $\gamma=0.1$, $\beta=0$
(d) $\gamma=0$, $\beta=2$
Figure 3: Energy behaviour for the FPU system with $p=1$, $T=100$, and $h_{2}=0.25$. Figure 3(a): the energy error computed by all schemes for the conservative system; Figure 3(b): the error of the true and deduced energy deviation by EKahan method shown in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy computed by all schemes for the two dissipative systems.
We test the convergence order and computational efficiency for all methods with those three settings of damping coefficients and the results are consistent. We report the results for the conservative equations in Figure 4. Figure 4(a) shows that Ekahan method is second-order and and Figure 4(b) shows that EKahan method is more efficient than all other methods. In addition, we observe that Kahan method is less accurate than any exponential integrator although it is shown to be more computationally effective than EAVF method. In Figure 5, we illustrate the numerical solutions given by EKahan method. Figure 5(a) shows that EKahan method captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time. Figure 5(b) shows that EKahan method captures the decay of the wave profiles as energy dissipates internally. Figure 3(c) and 5(b) demonstrate that although EKahan is designed for conservative equations, it can also capture the dissipative nature for the dissipative systems.
(a) order plot
(b) efficiency plot
Figure 4: Global error and computational time of different methods for the FPU system with $p=1$, $\gamma=0$, $\beta=0$, $T=100$ and different time step sizes $h_{i}$.
(a) $\gamma=0,\beta=0$
(b) $\gamma=0,\beta=2$
Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step need confirmation $h_{2}=0.25$.
FPU system with a fourth-order Hamiltonian
We consider $p=2$, i.e., the potential $U(y)$ is a quartic homogeneous polynomial function. We employ identical initial boundary condition and computational parameters to the $p=1$ case, with the exception of setting $\epsilon=100$. Figure 6(a) illustrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 6(b) verifies the formula of the energy deviation by EKahan method shown in Corollary 3.
Figure 6: Energy behaviour for the FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$. Left: the energy error computed by all schemes; right: the error of the true and deduced energy deviation computed by EKahan method shown in Corollary 3.
In Figure 7, we consider the global error and the computational cost using different time step sizes. Figure 7(a) demonstrates that EKahan method is second-order, and Figure 7(b) reveals its superior computational efficiency compared to all the other methods.
(a) order plot
(b) efficiency plot
Figure 7: Global error and computational time of different methods for the FPU system with $T=100$ and different time step sizes $h_{i}=1/2^{i}$, for $i=1,\dots,4$.
Zakharov–Kuznetsov equation
The Zakharov–Kuznetsov (ZK) equation is a two-dimensional generalization of the classical Korteweg–de Vries (KdV) equation [34], modeling the propagation of nonlinear waves in magnetized plasmas and stratified fluids. In this study, we consider the specific form
$u_{t}+u^{p}u_{x}+u_{xxx}+u_{xyy}=0,$
for $(x,y)\in[0,L]\times[0,L]$ and $t\in[0,T]$. Among its conserved quantities, we consider the following energy functional
$\mathcal{H}(t)=\frac{1}{2}\int_{\mathbb{R}^{2}}(u_{x}^{2}+u_{y}^{2})\,\mathrm{d}x\,\mathrm{d}y-\frac{1}{(p+1)(p+2)}\int_{\mathbb{R}^{2}}u^{p+2}\,\mathrm{d}x\,\mathrm{d}y.$
For $p=1$, the equation admits the following Hamiltonian structure:
$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$
To discretize the problem, we use the following discrete energy
$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u_{i,j})^{2}+(\delta_{x}^{-}u_{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u_{i,j})^{2}+(\delta_{y}^{-}u_{i,j})^{2}}{4}+\frac{u_{i,j}^{3}}{6}\right)\Delta x\Delta y.$
Here, $\delta_{x}^{+}$ and $\delta_{x}^{-}$, similarly $\delta_{y}^{+}$ and $\delta_{y}^{-}$, denote the forward and backward finite difference operators in the $x$- and $y$-directions, respectively. First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives, respectively. The corresponding discrete differential operators are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$. Using these notations, the semi-discrete system can be compactly written as
$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U.^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
where the vector $U\in\mathbb{R}^{(N_{x}-1)\times(N_{y}-1)}$ is flattened from $u_{i,j}$ by stacking columns successively, giving priority to the $y$-direction, and $U^{2}$ denotes the element-wise square of $U$. Here, $N_{x}$ and $N_{y}$ are the grid numbers in each spatial direction.
Numerical simulations are carried out on the spatial domain $[0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions. The spatial discretization employs a uniform grid with $N=33$ points in each spatial direction, resulting in a mesh size of $\Delta x=\Delta y=L/(N-1)$. For time integration, different time steps $h_{i}=0.01/2^{i+1},\;i=1,2,3,4$ are adopted. We consider the initial condition
$u(0,x,y)=\sqrt{2}[\sin(\frac{2\pi x}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi x}{L}+\frac{\pi}{4})]\times[\cos(\frac{2\pi y}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi y}{L}+\frac{\pi}{3})].$
Figure 8(a) demonstrates that the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 8(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 8: Energy behaviour for the Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$. Left: the energy error computed by EKahan, Kahan and LIEEP schemes; right: the error of the true and deduced energy deviation by EKahan shown in Theorem 2.
In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$. Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to Kahan method and LIEEP method.
(a) order plot
(b) efficiency plot
Figure 9: Global error and computational time of different methods for the Zakharov-Kuznetsov equation with $T=8$ and different time step sizes $h_{i}$.
Figure 10 illustrates the solution of the Zakharov-Kuznetsov equation. In Figure 10(a), we observe that the initial condition exhibits a regular, periodic pattern composed of multiple interacting modes. As time progresses to $t=8$, the solution evolves into a more localized, bell-shaped structure, reflecting the nonlinear dispersive nature of the equation; as shown in Figure 10(b). Despite the significant structural changes in the solution profile, the total energy of the system remains nearly constant throughout the simulation, demonstrating the excellent long-time stability.
Figure 10: Evolution to bell-shaped solitons in Zakharov-Kuznetsov equation, resolved with time step $h_{1}=0.0025$. Left: initial condition; right: soliton structure at $t=8$.
4 Conclusion
In this work, we proposed a linearly implicit, structure-preserving exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities by combining Kahan’s method with exponential integration techniques. The resulting scheme, EKahan method, successfully balances computational efficiency with the preservation of critical geometric structures, such as symmetry and near-conservation of energy. Compared to fully implicit exponential integrators, such as the EAVF method, EKahan requires only the solution of a single linear system per time step, significantly reducing computational costs. Although not all geometric properties of Kahan’s method are exactly inherited by EKahan method, we rigorously proved that EKahan maintains symmetry and a bounded and oscillatory energy error and achieves second-order convergence.
Through extensive numerical experiments on benchmark systems, we verified the method’s stability, long-term accuracy, and computational superiority. In particular, EKahan was shown to outperform not only the fully implicit scheme but also other linearly implicit methods, LIEEP and Kahan method, highlighting its enhanced efficiency among structure-preserving approaches. Furthermore, we provided the generalization of EKahan to problems involving high-order polynomial Hamiltonians, thereby broadening its applicability to a wider class of physical models.
Overall, the EKahan method offers a compelling tool for the long-time integration of semilinear Hamiltonian systems. Future work may explore extending this framework to adaptive step-size strategies, and constructing higher-order efficient structure-preserving exponential integrators.

### v2

@@
-A novel efficient structure-preserving exponential integrator for Hamiltonian systems
+arXiv:2506.07072v2[math.NA] 02 Mar 2026
 Abstract
-We propose a linearly implicit structure-preserving numerical method for semilinear Hamiltonian systems with polynomial nonlinearities, combining Kahan’s method and exponential integrator.
-This approach efficiently balances computational cost, accuracy and the preservation of key geometric properties, including symmetry and near-preservation of energy.
-By requiring only the solution of a single linear system per time step, the proposed method offers significant computational advantages while comparing with the state-of-the-art symmetric energy-preserving exponential integrators.
-The stability, efficiency and long-term accuracy of the method are demonstrated through numerical experiments on systems such as the Henon-Heiles system, the Fermi-Pasta-Ulam system and the two-dimensional Zakharov-Kuznestov equation.
+We develop an efficient linearly implicit time-reversible exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities.
+The method combines the variation of constants formula with Kahan-type discretization of the nonlinear term.
+For quadratic vector fields, we prove second-order convergence in finite dimensions and derive an exact stepwise energy-increment identity implying a precise notion of near conservation of energy.
+For higher-degree polynomial Hamiltonians, we employ polarization to obtain a linearly implicit multistep exponential extension and discuss its well-posedness, stability and convergence; we further provide a stepwise energy characterization that quantifies its energy drift in terms of the numerical increments.
+Numerical experiments on the Hénon-Heiles system, a continuum Fermi-Pasta-Ulam model, and the two-dimensional Zakharov-Kuznetsov equation corroborate the theoretical accuracy and demonstrate competitive accuracy-cost tradeoffs compared with representative symmetric energy-preserving exponential integrators.
 keywords:
-structure-preserving, energy preservation, linearly implicit , exponential integrator, symmetric
-1 Intorduction
+Hamiltonian system, linearly implicit, exponential integrato, symmetric, near conservation of energy
+1 Introduction
 This study is concerned with semilinear systems of the form:
@@
 $\begin{split}\dot{x}(t)&=Q\nabla H(x(t)),\\ x(t_{0})&=x_{0},\quad H(x(t))=\frac{1}{2}x^{\mathsf{T}}Mx+U(x),\end{split}$ (2)
-where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-order part [6].
+i.e., $A=QM$ and $f(x)=Q\nabla U(x)$, where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-degree part [5].
 For such systems, the Hamiltonian remains invariant along the exact solution trajectory.
@@
 Maintaining these geometric structures is essential for the long-time numerical integration of such systems, as it ensures accuracy and stability [15].
-To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 8], symplecticity-related exponential integrators [9, 10, 11, 12, 13], volume-preserving exponential integrator [14] and energy-preserving exponential integrators [15, 16].
+To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 22], symplecticity-related exponential integrators [34, 33, 26, 17, 30], volume-preserving exponential integrators [32] and energy-preserving exponential integrators [23, 12].
 Extensive studies have demonstrated that these specialized methods deliver significantly enhanced long-term performance compared to conventional high-order schemes.
@@
 This hybrid strategy has inspired the development of a broad class of exponential energy-preserving schemes.
-The computation of such fully implicit methods requires solving nonlinear systems at each time step—a process that can be computationally demanding.
+Solving such fully implicit methods requires tackling nonlinear systems at each time step, which can be computationally demanding.
 As an alternative, linearly implicit structure-preserving methods have been explored.
-These methods offer similar stability as fully explicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient.
-One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [24, 25, 18], which conserves a modified energy for semilinear Hamiltonian systems; however they are not symmetric.
+These methods offer similar stability to fully implicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient.
+One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [31, 24, 18], which conserves a modified energy for semilinear Hamiltonian systems but lacks symmetry.
+Symmetry is also an important geometric property, as it enforces discrete time-reversibility, yields an even-power error expansion, and often improves long-time behavior [15].
 A number of further developments along these lines can be found in [17, 30, 19, 35, 13].
 Another category involves multiple-point methods that portion out the nonlinearity [25, 22].
-These methods have the advantage of preserving both a modified energy and symmetry, but they are inherently multistep-even for systems with quadratic vector fields-which may bring stability concerns when compared to single-step method [28].
-In this work, we propose a linearly implicit structure-preserving method for semilinear Hamiltonian systems with polynomial nonlinearities, by combining Kahan’s method [29] with exponential integrators.
-Kahan’s method is employed due to several favorable properties: i) it is linearly implicit, requiring only the solution of a single linear system per time step, which improves computational efficiency compared to fully implicit methods; ii) it is symmetric, a key property for ensuring stable long-term integration; and iii) it preserves a modified energy and measure, geometric characteristics that are also satisfied by the exact flow of the original system.
-The proposed method inherits all advantageous properties of Kahan’s method except for the exact conservation of a modified energy and measure.
-However, we show that the energy error remains bounded and exhibits oscillatory behavior over time.
-Moreover, we derive an explicit analytical expression for the energy error, offering insights into the method’s long-term energy behavior.
-Through numerical experiments, we demonstrate that the proposed method achieves superior computational efficiency compared to both the fully implicit exponential AVF method and the linearly implicit method based on the polarized discrete gradient.
-Additionally, we provide a rigorous proof that the proposed method achieves second-order convergence.
+These methods have the advantage of preserving both a modified energy and symmetry, but they may suffer from stability concerns when compared to Kahan methods [9], and a systematic convergence theory for linearly implicit time-reversible exponential integrators, especially in the higher-degree polynomial setting, is still limited [22].
+These observations motivate the design of stable time-reversible and linearly implicit exponential integrators, together with a rigorous convergence analysis.
+We therefore incorporate the Kahan discretization for the nonlinear term, since it is linearly implicit, time-reversible and known to preserve a modified energy and a modified measure [6, 7].
+To treat the stiff linear dynamics in (1), we embed Kahan update into the variation of constants exponential framework, leading to symmetric schemes that require only a linear solve per step and admit a quantitative near conservation of energy.
+This coupling is, however, nontrivial: the exponential update changes the algebraic structure of the original Kahan discretization, so its classical modified-energy preservation mechanism does not carry over in general, and the energy conservation analysis must be rebuilt.
+Our main contributions in this work are as follows.
+First, for quadratic vector fields (cubic Hamiltonians), we obtain a one-step linearly implicit symmetric exponential scheme and prove second-order convergence on finite time intervals.
+Second, for higher-degree polynomial Hamiltonians, we derive a linearly implicit time-reversible multistep exponential extension via polarization and establish solvability and stability analysis.
+Third, we derive explicit stepwise energy increment identities, which yield computable energy-drift estimates in terms of numerical increments and explain the favorable long-time behavior observed in our experiments.
 The remainder of this paper is organized as follows.
-In Section 2, we introduce the proposed linearly implicit structure-preserving exponential integrator and discuss its key properties, including symmetry, energy preservation and convergence.
+In Section 2, we introduce the proposed linearly implicit exponential integrator for cubic Hamiltonians, discuss its key properties, including symmetry, convergence, near preservation of energy, and its generalization to systems with higher-degree polynomial Hamiltonians.
 In Section 3, we present numerical experiments to demonstrate the stability, computational efficiency, and long-term behavior of the proposed method using different types of differential equations.
@@
 2 The new linearly implicit structure-preserving symmetric method
-Before proposing the new exponential integrator, we will introduce Kahan’s method. For a Hamiltonian system with a cubic Hamiltonian, Kahan’s method has the form
+Before proposing the new exponential integrator, we will introduce Kahan’s method and exponential integrators since our new scheme is created based on them.
+For system (2) with a cubic Hamiltonian, Kahan’s method has the form
 $x_{n+1}=Q(-\frac{1}{2}\nabla H(x_{n})+2\nabla H(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla H(x_{n+1})),$
@@
 These properties guarantee the superior numerical behavior of Kahan’s method during long-time integration [9].
-For a semilinear system (1), we consider the following variation of constants formula on interval $[t_{0},t_{0}+h]$
-$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$ (3)
-The key idea of constructing exponential integrator is to integrate the linear part of equation (3) exactly meanwhile finding an approximation to the integration on the right-hand side.
+For a semilinear system (1), we consider the following variation of constants formula on interval $t_{0},t_{0}+h]$
+$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$
+The key idea of constructing exponential integrator is to integrate the linear part exactly meanwhile finding an approximation to the integration of the nonlinear part.
 For example, the exponential Euler method is obtained by approximating the nonlinear term by $f(x(t_{0}))$ [16], thereby having the form
 $x_{n+1}=e^{hA}x_{n}+h\phi(hA)f(x_{n}),$
-where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field method is obtained by approximating the nonlinear term by the discrete gradient [15], possessing the form
-[…]
-In this work, we approximate the nonlinear term using Kahan’s method, leading to the new linearly implicit exponential integrator suitable for solving semilinear systems (1):
+where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field (EAVF) method is obtained by approximating the nonlinear term by the discrete gradient [23], possessing the form
+$x_{n+1}=e^{hA}x_{n}+h\phi(hA)\int_{0}^{1}f(\xi x_{n}+(1-\xi)x_{n+1})d\xi.$
+2.1 The one-step EKahan scheme for cubic Hamiltonians
+For system (2) with a cubic Hamiltonian, we propose a linearly implicit exponential integrator by approximating the nonlinear term using the above mentioned Kahan discretization, denoted by EKahan
 $x_{n+1}=e^{hA}x_{n}+h\phi(hA)(-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})).$ (3)
-We denote this newly derived scheme as the EKahan scheme, and its properties are described in detail below.
+Before presenting the properties of scheme (3), we give the following notations and assumptions:
+Norms and growth constants
+Let $\|\cdot\|$ be an induced matrix norm and the associated vector norm. For a fixed final time $T>0$, define
+$M_{E}=\max_{0\leq t\leq T}\|e^{tA}\|,\qquad M_{\phi}=\max_{0\leq t\leq T}\|\phi(tA)\|.$
+Lipschitz assumptions
+For system (2), we assume that there exists a convex bounded set $\mathcal{D}\subset\mathbb{R}^{d}$ such that both the exact solution $\{x(t):t\in[0,T]\}$ and the numerical solution $\{x_{n}:0\leq nh\leq T\}$ remain in $\mathcal{D}$, and that $f$ is Lipschitz on $\mathcal{D}$ with constant $L$.
 Lemma 1.
-The EKahan scheme (4) is symmetric.
+Assuming that $f$ is a quadratic vector field and that the exact solution of (1) is sufficiently smooth on $t_{n},t_{n+1}]$, then the EKahan scheme (3) has local truncation error $O(h^{3})$.
 Proof.
-[…]
-$x_{n}=e^{-Ah}x_{n+1}-h\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (5)
-Multiplying both sides of (5) by $e^{hA}$, we get
-$e^{hA}x_{n}=x_{n+1}-he^{hA}\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (6)
-Using the identiy $\phi(-V)=e^{-V}\phi(V)$ for any matrix $V$, we can rewrite the above equation (6) as
-$e^{hA}x_{n}=x_{n+1}-h\phi(hA)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$
-This equation is equivalent to (4), confirming that the EKahan scheme (4) is symmetric. ∎
-Lemma 2.
-The EKahan scheme (4) is second-order.
+For quadratic $f$, we have the exact Taylor formulas
+$f(x_{n+1})$ $=f(x_{n})+f^{\prime}(x_{n})\Delta_{x}+\tfrac{1}{2}f^{\prime\prime}(x_{n})(\Delta_{x},\Delta_{x}),$
+$f\left(\tfrac{x_{n}+x_{n+1}}{2}\right)$ $=f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}+\tfrac{1}{8}f^{\prime\prime}(x_{n})(\Delta_{x},\Delta_{x}),$
+where $\Delta_{x}=x_{n+1}-x_{n}$ and $f^{\prime\prime}$ is constant. Hence, we have
+$\tfrac{1}{2}f(x_{n})+2f\!\left(\tfrac{x_{n}+x_{n+1}}{2}\right)-\tfrac{1}{2}f(x_{n+1})=f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}.$ (4)
+Using (4), the EKahan scheme (3) can be rewritten as
+$x_{n+1}=e^{hA}x_{n}+h\phi(hA)\Bigl(f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}\Bigr).$ (5)
+Since $e^{hA}=I+h\phi(hA)A$ and $A$ commutes with $\phi(hA)$, (5) becomes
+$x_{n+1}=x_{n}+h\phi(hA)g(x_{n})+\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Delta_{x},\quad\text{with}\quad g(x)=Ax+f(x),$
+and therefore we have
+$\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)\Delta_{x}=h\phi(hA)g(x_{n}).$ (6)
+For sufficiently small $h$, the matrix in (6) is invertible, and we may write
+$\Delta_{x}=h\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)^{-1}\phi(hA)g(x_{n}).$ (7)
+Using the series expansion $\phi(hA)=I+\tfrac{h}{2}A+O(h^{2})$ and the Neumann expansion
+$\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)^{-1}=I+\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})+O(h^{2}),$
+we obtain from (7) that
+$\Delta_{x}=h\,g(x_{n})+\tfrac{h^{2}}{2}\bigl(A+f^{\prime}(x_{n})\bigr)g(x_{n})+O(h^{3}).$
+Consequently, we have
+$x_{n+1}=x_{n}+h\,g(x_{n})+\tfrac{h^{2}}{2}\bigl(A+f^{\prime}(x_{n})\bigr)g(x_{n})+O(h^{3}).$ (8)
+On the other hand, the exact solution satisfies
+$x(t_{n+1})=x(t_{n})+h\,g(x(t_{n}))+\tfrac{h^{2}}{2}g^{\prime}(x(t_{n}))g(x(t_{n}))+O(h^{3}),$ (9)
+where $g^{\prime}(x)=A+f^{\prime}(x)$. Taking $x_{n}=x(t_{n})$ and comparing (8) with (9) yields $x_{n+1}-x(t_{n+1})=O(h^{3})$.
+Theorem 1.
+Assume $f$ satisfies the standard Lipschitz condition and the local truncation error satisfies $\|\tau_{n+1}\|\leq C_{\rm loc}h^{3}$ for all $0\leq nh\leq T$, which is true by Lemma 1.
+Then there exist constants $C_{1},C_{2}>0$, depending only on $\|A\|$, $L$, $T$, $M_{E}$, $M_{\phi}$ and $C_{\rm loc}$, such that
+$\|x(t_{n})-x_{n}\|\leq e^{C_{1}T}\Bigl(\|x(t_{0})-x_{0}\|+C_{2}h^{2}\Bigr),\quad 0\leq nh\leq T,$
+if the step size satisfies
+$h\leq h_{0}:=\frac{1}{3LM_{\phi}}.$ (10)
+In particular, if $x_{0}=x(t_{0})$, then $\max\limits_{0\leq nh\leq T}\|x(t_{n})-x_{n}\|=O(h^{2})$.
 Proof.
-By applying Taylor expansion to $f(\frac{x_{n}+x_{n+1}}{2})$ and $f(x_{n+1})$ at the point $x_{n}$ and using the fact that $f$ is quadratic, we can obtain
-$-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})=f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n}).$
-Therefore, EKahan scheme (4) can be rewritten as
-$x_{n+1}$ $=e^{hA}x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})].$ (7)
-Since the matrix exponential satisfies $e^{hA}=I+h\phi(hA)A$, we rewrite equation (7) as
-$x_{n+1}$ $=(I+h\phi(hA)A)x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})]$
-$=x_{n}+h\phi(hA)g(x_{n})+\frac{h}{2}\phi(hA)f^{\prime}(x_{n})(x_{n+1}-x_{n}),$ (8)
-where the function $g(x)$ is the right-hand side of the semilinear equation (1), i.e., $g(x):=Ax+f(x)$.
-Supposing the matrix $I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n})$ is invertible, solving for $x_{n+1}-x_{n}$ from equation (2) gives
-$x_{n+1}-x_{n}=h(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n}).$
-By $L^{\prime}H\hat{o}pital^{\prime}s\;rule$, we have $\lim\limits_{z\rightarrow 0}\frac{e^{z}-1}{z}=1$, so $\lim\limits_{h\rightarrow 0}\phi(hA)=I$ and
-$\lim\limits_{h\rightarrow 0}(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n})=g(x_{n}).$
-Therefore, we have the following truncation for the EKahan scheme (7)
-$x_{n+1}$ $=x_{n}+hg(x_{n})+O(h^{2}).$ (9)
-By Taylor expansion, the exact solution of the semilinear Hamiltonian system (1) satisfies
-$x(t_{n+1})=x(t_{n})+hg(x(t_{n}))+\frac{h^{2}}{2}(A+f^{\prime}(x(t_{n})))g(x(t_{n}))+O(h^{3}).$ (10)
-Supposing $x_{n}$ is exactly solved and subtracting equation (10) from equation (9), we get
-$x_{n+1}-x(t_{n+1})=O(h^{2}).$
-Moreover, EKahan scheme (4) is symmetric from Lemma 1, confirming that this method is at least second-order. ∎
-Theorem 1.
-[…]
-$\|f(x_{2})-f(x_{1})\|\leq L\|x_{2}-x_{1}\|.$
-Then, the error bound
-$\|x(t_{n})-x_{n}\|\leq e^{k_{2}T}(k_{1}\|x(t_{0})-x_{0}\|+k_{3}h^{2})$
-holds in $0\leq nh\leq T$ for constant $k_{1},\;k_{2},\;k_{3}$.
-Proof.
-The assumptions about the operator $A$ guarantee the boundness of $\|\phi(hA)\|$ and $\|e^{jhA}\|$, $j=1,2,\dots$, and we denote the bound by $M$; see Lemma 2.4 from [5].
-Since EKahan scheme (4) is second-order, we get the following truncation when inserting the exact solutions $x(t_{n})$ and $x(t_{n+1})$ to the numerical scheme:
-$\begin{split}x(t_{n+1})=&e^{hA}x(t_{n})+h\phi(hA)[-\frac{1}{2}f(x(t_{n}))-\frac{1}{2}f(x(t_{n+1}))\\ &+2f(\frac{x(t_{n})+x(t_{n+1})}{2})]+O(h^{3}).\end{split}$ (11)
-Denote the numerical error terms by $\varepsilon_{n}=x(t_{n})-x_{n}$ and
-$\delta_{n+1,n}=$ $-\frac{1}{2}f(x(t_{n}))+2f(\frac{x(t_{n})+x(t_{n+1})}{2})-\frac{1}{2}f(x(t_{n+1}))$
-$-\big{(}-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})\big{)}.$
-Subtracting equation (4) from equation(11), we get the following error recursion equation
-$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+O(h^{3}).$
-This recursion equation gives
-$\varepsilon_{n}=e^{nAh}\varepsilon_{0}+\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}+\sum\limits_{j=0}^{n-1}e^{jAh}O(h^{3}).$ (12)
-Since $f$ satisfies $Lipschitz\;condition$, we have
-$\|\delta_{n-j,n-j-1}\|\leq\frac{3}{2}L\|\varepsilon_{n-j}\|+\frac{3}{2}L\|\varepsilon_{n-j-1}\|.$
-Then we have
-$\|\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}\|\leq\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|.$ (13)
-Combing (12) and (13), we obtain
-$\|\varepsilon_{n}\|\leq\|e^{TA}\|\|\varepsilon_{0}\|+\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+nMO(h^{3}).$ (14)
-For small $h$ such that $1-\frac{3}{2}hLM$ is positive, suppose there exists a positive real number $k_{3}$ satisfying $\frac{nMO(h^{3})}{1-\frac{3}{2}hLM}\leq k_{3}h^{2}$.
-Denote by $k_{1}=\frac{\|e^{TA}\|\|\varepsilon_{0}\|}{1-\frac{3}{2}hLM}$ and $k_{2}=\frac{3LM^{2}}{1-\frac{3}{2}hLM}$.
-Then we can obtain the following estimation for the numerical error $\varepsilon_{n}$ from the above relation (14):
-$\|\varepsilon_{n}\|$ $\leq k_{1}\|\varepsilon_{0}\|+k_{2}h\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+k_{3}h^{2}$
-$\leq e^{nk_{2}h}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2})$
-$=e^{k_{2}T}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2}).$
-The last inequality follows from the Gronwall’s inequality. ∎
-Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variational constant formula applied to the semilinear system (2).
+Set $\varepsilon_{n}=x(t_{n})-x_{n}$ and $\Psi(u,v)=-\tfrac{1}{2}f(u)+2f\!\Bigl(\tfrac{u+v}{2}\Bigr)-\tfrac{1}{2}f(v)$. Inserting the exact solution into (3) yields
+$x(t_{n+1})=e^{hA}x(t_{n})+h\phi(hA)\Psi(x(t_{n}),x(t_{n+1}))+\tau_{n+1},\quad\|\tau_{n+1}\|\leq C_{\rm loc}h^{3}.$ (11)
+Subtracting (3) from (11) gives the error recursion
+$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+\tau_{n+1},$ (12)
+where $\delta_{n+1,n}=\Psi(x(t_{n}),x(t_{n+1}))-\Psi(x_{n},x_{n+1})$. Since $\mathcal{D}$ is convex and $f$ is Lipschitz on $\mathcal{D}$, we have
+$\|\delta_{n+1,n}\|$ $\leq\tfrac{1}{2}\|f(x(t_{n}))-f(x_{n})\|+2\Bigl\|f\!\Bigl(\tfrac{x(t_{n})+x(t_{n+1})}{2}\Bigr)-f\!\Bigl(\tfrac{x_{n}+x_{n+1}}{2}\Bigr)\Bigr\|$
+$+\tfrac{1}{2}\|f(x(t_{n+1}))-f(x_{n+1})\|$
+$\leq\tfrac{3}{2}L(\|\varepsilon_{n}\|+\|\varepsilon_{n+1}\|).$
+Taking norms in (12) and using $\|\phi(hA)\|\leq M_{\phi}$ yields
+$\|\varepsilon_{n+1}\|\leq\|e^{hA}\|\,\|\varepsilon_{n}\|+\tfrac{3}{2}LhM_{\phi}\bigl(\|\varepsilon_{n}\|+\|\varepsilon_{n+1}\|\bigr)+C_{\rm loc}h^{3}.$ (13)
+Denoting by $\alpha=\tfrac{3}{2}LhM_{\phi}$, moving the term $\alpha\|\varepsilon_{n+1}\|$ to the left in (13) gives
+$(1-\alpha)\|\varepsilon_{n+1}\|\leq(\|e^{hA}\|+\alpha)\|\varepsilon_{n}\|+C_{\rm loc}h^{3},$
+and hence we have
+$\|\varepsilon_{n+1}\|\leq q(h)\,\|\varepsilon_{n}\|+\frac{C_{\rm loc}}{1-\alpha}h^{3},\quad\text{with}\quad q(h)=\frac{\|e^{hA}\|+\alpha}{1-\alpha}.$ (14)
+Using $e^{hA}=I+\int_{0}^{h}Ae^{sA}\mathrm{d}s$ and $\|e^{sA}\|\leq M_{E}$ for $0\leq s\leq h$, we obtain
+$\|e^{hA}\|\leq 1+h\|A\|M_{E}.$
+Moreover, under (10) we have $\alpha\leq\tfrac{1}{2}$, so that $(1-\alpha)^{-1}\leq 1+2\alpha$ and $2\alpha^{2}\leq\alpha$. Therefore, we get
+$q(h)=\frac{\|e^{hA}\|+\alpha}{1-\alpha}\leq(\|e^{hA}\|+\alpha)(1+2\alpha)\leq 1+Ch$
+with $C=\|A\|M_{E}+3(1+M_{E})LM_{\phi}$. In addition, since $\alpha\leq\tfrac{1}{2}$ we have $(1-\alpha)^{-1}\leq 2$, and thus
+$\frac{C_{\rm loc}}{1-\alpha}h^{3}\leq 2C_{\rm loc}h^{3}.$
+Substituting these bounds into (14) gives
+$\|\varepsilon_{n+1}\|\leq(1+Ch)\|\varepsilon_{n}\|+2C_{\rm loc}h^{3},\qquad 0\leq nh\leq T,\ \ h\leq h_{0}.$ (15)
+Applying the discrete Gronwall inequality to (15) yields, for $t_{n}=nh\leq T$,
+$\|\varepsilon_{n}\|\leq(1+Ch)^{n}\|\varepsilon_{0}\|+2C_{\rm loc}h^{3}\sum_{j=0}^{n-1}(1+Ch)^{j}.$
+Using $(1+Ch)^{n}\leq e^{CT}$ for $nh\leq T$ and
+$\sum_{j=0}^{n-1}(1+Ch)^{j}=\frac{(1+Ch)^{n}-1}{Ch}\leq\frac{(1+Ch)^{n}}{Ch}\leq\frac{e^{CT}}{Ch},$
+we obtain
+$\|\varepsilon_{n}\|\leq e^{CT}\|\varepsilon_{0}\|+\frac{2C_{\rm loc}}{C}e^{CT}h^{2},$
+which proves the stated estimate with $C_{1}=C$ and $C_{2}=2C_{\rm loc}/C$. ∎
+Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variation of constant formula applied to the semilinear system (2).
 The proof of this lemma is inspired by Theorem 1 in [23].
 Lemma 2.
-For the semilinear system (2), any exponential integrator with the form
+For system (2), any exponential integrator of the form
 $x_{n+1}=e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1}),$
 satisfies the following equation
-$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n+1},x_{n})=0,$
-where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
-Proof.
-Assuming that the matrix $M$ is not singular and denoting by $V=hA=hQM$ and ^$\hat{\nabla}U=M\breve{\nabla}U$, we get the following equation after replacing $x_{n+1}$ by $e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1})$:
-$\begin{split}&\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}{x_{n}}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1})\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+h{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+\frac{1}{2}h^{2}\hat{\nabla}U^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}-I)\hat{\nabla}U+{\hat{\nabla}U}^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}\hat{\nabla}U\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}(e^{V}-I)^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{\breve{\nabla}U}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U\\ &=\frac{1}{2}(x_{n}+\breve{\nabla}U)^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)(x_{n}+\breve{\nabla}U)+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-Me^{V})\breve{\nabla}U\\ &=0.\end{split}$
-The last equation follows from the fact that ${e^{V}}^{\mathsf{T}}Me^{V}-M$ and ${e^{V}}^{\mathsf{T}}M-Me^{V}$ are skew symmetric.
-If $M$ is singular, we follow the idea in [15], introducing a series of symmetric and nonsingular matrices $M_{\delta}$ which converges to $M$ when $\delta$ approaches zero to complete the proof. ∎
-The discretized gradient ^$\hat{\nabla}U(x_{n},x_{n+1})$ can include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method.
-We denote the discretized gradient obtained via Kahan’s method by $\nabla U_{\text{K}}$, which can be written as
-$\nabla U_{\text{K}}=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
+$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n},x_{n+1})=0,$
+where $\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
+Proofs of this lemma and the subsequent lemmas and corollaries are given in the Appendix.
+The discretized gradient $\hat{\nabla}U(x_{n},x_{n+1})$ include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method.
+We denote the discretized gradient obtained via Kahan’s method by $\nabla_{\mathrm{K}}U$, which can be written as
+$\nabla_{\mathrm{K}}U=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
+For the AVF choice, $\hat{\nabla}U$ is a discrete gradient and Lemma 2 implies exact energy conservation. For the Kahan choice, Lemma 2 will yield the following energy-increment identity.
 Theorem 2.
-For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, EKahan scheme (4) nearly preserves the exact energy and has the following error estimation
-$H_{n+1}-H_{n}$ $=U(x_{n+1}-x_{n}),$
+For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, the EKahan scheme (3) satisfies the stepwise identity
+$H_{n+1}-H_{n}=U(x_{n+1}-x_{n}),$ (16)
 where $H_{n}=\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n})$ is the discrete energy at $t=t_{n}$.
 Proof.
-Based on Lemma 3, we have
-$\begin{split}H_{n+1}-H_{n}&=\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n+1})-U(x_{n})\\ &=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).\end{split}$ (15)
-For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla U_{\text{K}}=\frac{U^{\prime\prime}(x_{n+1})x_{n}}{2}$.
-Then, the above energy error (15) can be reformulated as
+Applying Lemma 2 with $\hat{\nabla}U=\nabla_{\mathrm{K}}U$ gives
+$H_{n+1}-H_{n}=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).$ (17)
+For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla_{\mathrm{K}}U=\frac{1}{2}U^{\prime\prime}(x_{n+1})x_{n}$. Then (17) can be reformulated as
 $H_{n+1}-H_{n}$
-$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
-$-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
-$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
-$+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
-$-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
+$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
+$+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
+$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
+$-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
 $=\frac{1}{6}(x_{n+1}-x_{n})^{\mathsf{T}}(U^{\prime\prime}(x_{n+1})x_{n+1}+U^{\prime\prime}(x_{n})x_{n}-2U^{\prime\prime}(x_{n+1})x_{n})$
@@
 $=U(x_{n+1}-x_{n}).$
-For Hamiltonian systems (2) with high-order polynomial function $U(x)$, EKahan method can be generalized using the polarization technique presented in [30].
-Suppose $U(x)$ is a homogeneous polynomial of order $k+2$, EKahan scheme is given by
-$x_{n+k}=e^{khQM}x_{n}+kh\phi(khQM)Q\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k}),$ (16)
-where $\nabla U_{\text{K}}(\cdot,\cdots,\cdot)$ is a symmetric $k+1$-linear polarized function of $\nabla U(x)$, satisfying
-$\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$
-In this case, we give the following corollaries to describe EKahan’s properties.
 Corollary 1.
-For semilinear system (2) with high-order polynomial function $U(x)$, the EKahan scheme (16) is symmetric and second order if the initial starting points $x_{n+1},\cdots,x_{n+k}$ are given using a at least second-order method.
-This corollary is a generalization of Lemma 1 and 2 for Hamiltonian systems (2) with cubic polynomial function $U(x)$ to systems with high-order polynomial functions, and the proof follows similarly.
+Let the potential $U$ be a general cubic polynomial with homogeneous decomposition as
+$U(x)=U_{3}(x)+U_{2}(x)+U_{1}(x)+U_{0},$
+where $U_{j}$ is homogeneous of degree $j$. Then the EKahan scheme (3) satisfies
+$H_{n+1}-H_{n}=U_{3}(x_{n+1}-x_{n}),$
+showing that the stepwise energy increment depends only on the cubic component.
+Remark 1.
+Theorem 2 and Corollary 1 provide a computable description of the one-step energy defect in terms of the numerical increment.
+This complements standard finite-time estimates obtained indirectly from second-order convergence, and explains the favourable long-time behaviour of EKahan despite not implying exact energy conservation or symplecticity for general nonlinear Hamiltonian systems.
+2.2 The multistep EKahan scheme for higher-degree polynomial Hamiltonians
+We now extend the one-step EKahan scheme (3) from cubic potentials to polynomial potentials of degree $k+2$ with $k\in\mathbb{Z}^{+}$.
+The key point is to retain the two features of EKahan: (i) exponential propagation of the linear flow, and (ii) a Kahan-type reversible discretization of the polynomial Hamiltonian part that requires solving only one linear system per time step.
+To achieve (ii), we portion the nonlinearity by employing the symmetric $(k+1)$-linear polarization introduced in [6], leading to a $k$-step method.
+We denote the symmetric $(k+1)$-linear polarization of $\nabla U$ by
+$\nabla_{\mathrm{K}}U(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\,\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$ (18)
+Then, we propose the following $k$-step EKahan scheme
+$x_{n+k}=e^{khA}x_{n}+kh\phi(khA)Q\nabla_{\mathrm{K}}U(x_{n},x_{n+1},\cdots,x_{n+k}).$ (19)
+Lemma 3.
+Let $U$ be a polynomial of degree $k+2$.
+Then for each $n$ there exist a matrix $B_{n}\in\mathbb{R}^{d\times d}$ and a vector $\ell_{n}\in\mathbb{R}^{d}$, depending only on $(x_{n},\dots,x_{n+k-1})$, such that
+$\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k})=B_{n}x_{n+k}+\ell_{n}.$ (20)
+Consequently, (19) is equivalent to the linear system
+$\Bigl(I-kh\,\phi(khQM)\,QB_{n}\Bigr)x_{n+k}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n}.$
+In particular, if $U$ is homogeneous, then $\ell_{n}=0$.
+We list several physically relevant polynomial potentials and give the corresponding polarized gradient $\nabla_{\mathrm{K}}U$.
+  • For the standard lattice Duffing oscillator with on-site potential
+    $U(q)=\frac{\beta}{4}\sum_{j=1}^{d}(q^{j})^{4},\qquad\beta>0,$
+    the polarization formula (18) (with $k=2$) yields
+    $\nabla_{\mathrm{K}}U(q_{n},q_{n+1},q_{n+2})=\beta\,(q_{n}\odot q_{n+1}\odot q_{n+2}),$
+    where $\odot$ denotes componentwise product, and the corresponding coefficient matrix and vector in (20) are $B_{n}=\beta\,\mathrm{diag}(q_{n}\odot q_{n+1})$ and $\ell_{n}=0$.
+  • For the Fermi-Pasta-Ulam with quartic potential
+    $U(q)=\frac{\beta}{4}\sum_{i=1}^{N-1}\bigl(r^{i}(q)\bigr)^{4},\qquad r^{i}(q)=\frac{q^{i+1}-q^{i}}{\Delta x},$
+    introduce the strain vector $r(q)=(r^{1}(q),\ldots,r^{N-1}(q))^{\mathsf{T}}$ and $r_{n}=r(q_{n})$. Then polarization formula (18) (with $k=2$) gives
+    $\nabla_{\mathrm{K}}U(q_{n},q_{n+1},q_{n+2})=D^{\mathsf{T}}\bigl(\beta\,(r_{n}\odot r_{n+1}\odot r_{n+2})\bigr),$
+    where $D$ is the forward difference operator, and the corresponding coefficient matrix and vector in (20) are $B_{n}=D^{\mathsf{T}}\,\mathrm{diag}(w_{n})\,D$ and $\ell_{n}=0$, with $w_{n}=\beta\,(r_{n}\odot r_{n+1})$.
+  • For the Hénon-Heiles potential
+    $U(p,q)=p^{2}q-\frac{1}{3}q^{3},$
+    the two-point Kahan discretized gradient (with $k=1$) is
+    $\nabla_{\mathrm{K}}U(r_{n},r_{n+1})=\begin{pmatrix}p_{n}q_{n+1}+p_{n+1}q_{n}\\ p_{n}p_{n+1}-q_{n}q_{n+1}\end{pmatrix}\qquad r=(p,q)^{\mathsf{T}},$
+    and the corresponding coefficient matrix and vector in (20) are $B_{n}=\begin{pmatrix}q_{n}&p_{n}\\ p_{n}&-q_{n}\end{pmatrix}$ and $\ell_{n}=0$.
+Before presenting the properties of scheme (19), we give the following notations and assumptions:
+Stacked state and history norm
+Define the stacked state
+$Y_{n}=(x_{n+k-1},x_{n+k-2},\dots,x_{n})\in(\mathbb{R}^{d})^{k},$
+and the norm on the $k$-step history
+$\|Y_{n}\|_{\infty}=\max_{0\leq j\leq k-1}\|x_{n+j}\|.$
+$k$-step assumptions
+For the $k$-step scheme (19), we assume that along the numerical trajectory the matrices $B_{n}$ in Lemma 3 satisfy $\|B_{n}\|\leq M_{B}$, and the uniform bound $M_{\phi}=\max\limits_{0\leq s\leq kT}\|\phi(sA)\|$ holds for a fix final time $T>0$.
+Finally, we assume the mappings $(x_{n},\ldots,x_{n+k-1})\mapsto B_{n}$ and $(x_{n},\ldots,x_{n+k-1})\mapsto\ell_{n}$ are Lipschitz on the relevant bounded history set.
+Theorem 3.
+Assuming the $k$-step assumptions, then the $k$-step EKahan scheme (19) is well-defined for a sufficiently small $h$ and stable on $[0,T]$ with respect to starting perturbations.
+Proof.
+By Lemma 3, (19) is equivalent to
+$\bigl(I-kh\,\phi(khQM)\,QB_{n}\bigr)x_{n+k}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n}.$ (21)
+Under the $k$-step assumptions, there will exist a $h_{0}$ such that
+$kh\,M_{\phi}\,\|Q\|\,M_{B}\leq\theta<1\quad\text{for all }0<h\leq h_{0}.$ (22)
+Then the coefficient matrix $I-kh\,\phi(khA)\,QB_{n}$ is invertible and Neumann-series bound further gives
+$\bigl\|\bigl(I-kh\,\phi(khA)\,QB_{n}\bigr)^{-1}\bigr\|\leq\frac{1}{1-\theta}\quad\text{for all }n\text{ and }0<h\leq h_{0}.$ (23)
+Therefore $x_{n+k}$ is uniquely determined.
+Let $\{x_{n}\}$ and $\{\tilde{x}_{n}\}$ be two numerical solutions produced by (19) with the same step size.
+Set the error of the current state by $\delta_{n}=x_{n}-\tilde{x}_{n}$ and the error of the stacked state by $\Delta_{n}=Y_{n}-\tilde{Y}_{n}$ .
+Define
+$G_{n}=kh\,\phi(khQM)\,QB_{n},\qquad b_{n}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n},$
+and $\tilde{G}_{n},\tilde{b}_{n}$ analogously. Then from (21) we have $(I-G_{n})x_{n+k}=b_{n}$ and $(I-\tilde{G}_{n})\tilde{x}_{n+k}=\tilde{b}_{n}$. Subtracting and using the resolvent identity yields
+$\delta_{n+k}=(I-G_{n})^{-1}(b_{n}-\tilde{b}_{n})+(I-G_{n})^{-1}(G_{n}-\tilde{G}_{n})(I-\tilde{G}_{n})^{-1}\tilde{b}_{n}.$ (24)
+Taking $h_{0}$ such that $\theta=\frac{1}{2}$ in (22)-(23), we have $kh\,M_{\phi}\,\|Q\|\,M_{B}\leq\tfrac{1}{2}$ and
+$\|(I-G_{n})^{-1}\|\leq\frac{1}{1-\|G_{n}\|}\leq 1+2\|G_{n}\|\leq 1+C_{B}h,\quad C_{B}:=2kM_{\phi}\|Q\|M_{B},$ (25)
+and the same bound holds for $\|(I-\tilde{G}_{n})^{-1}\|$.
+Moreover, recalling from the proof of Theorem 1, we have $\|e^{khQM}\|\leq 1+C_{E}h$ with $C_{E}=k\|QM\|M_{E}$. Taking norms in (24) and using (25) yield
+$\|\delta_{n+k}\|\leq(1+C_{B}h)\,\|b_{n}-\tilde{b}_{n}\|+(1+C_{B}h)^{2}\,\|G_{n}-\tilde{G}_{n}\|\,\|\tilde{b}_{n}\|.$ (26)
+Furthermore, we have
+$\begin{split}\|b_{n}-\tilde{b}_{n}\|&\leq\|e^{khQM}\|\,\|\delta_{n}\|+kh\,\|\phi(khQM)\|\,\|Q\|\,\|\ell_{n}-\tilde{\ell}_{n}\|\\ &\leq(1+C_{E}h)\|\delta_{n}\|+kh\,M_{\phi}\|Q\|\,\|\ell_{n}-\tilde{\ell}_{n}\|,\end{split}$
+and
+$\|G_{n}-\tilde{G}_{n}\|\leq kh\,\|\phi(khQM)\|\,\|Q\|\,\|B_{n}-\tilde{B}_{n}\|\leq kh\,M_{\phi}\|Q\|\,\|B_{n}-\tilde{B}_{n}\|.$
+By the Lipschitz assumption on the mappings $(x_{n},\ldots,x_{n+k-1})\mapsto B_{n}$ and $(x_{n},\ldots,x_{n+k-1})\mapsto\ell_{n}$, there exist constants $L_{B},L_{\ell}>0$ such that
+$\|B_{n}-\tilde{B}_{n}\|\leq L_{B}\|\Delta_{n}\|_{\infty},\qquad\|\ell_{n}-\tilde{\ell}_{n}\|\leq L_{\ell}\|\Delta_{n}\|_{\infty}.$
+In addition, boundedness of the trajectories implies $\|\tilde{b}_{n}\|\leq M_{b}$ for some constant $M_{b}>0$.
+Substituting these estimates into (26) and using $\|\delta_{n}\|\leq\|\Delta_{n}\|_{\infty}$ gives
+$\|\delta_{n+k}\|\leq\Bigl((1+C_{B}h)(1+C_{E}h+khM_{\phi}\|Q\|L_{\ell})+kh(1+C_{B}h)^{2}M_{\phi}\|Q\|L_{B}\,M_{b}\Bigr)\|\Delta_{n}\|_{\infty}.$
+Therefore, for $0<h\leq h_{0}$ there exists $\widetilde{C}>0$ independent of $h$ such that
+$\|\delta_{n+k}\|\leq(1+\widetilde{C}h)\|\Delta_{n}\|_{\infty}.$
+Consequently,
+$\|\Delta_{n+1}\|_{\infty}=\max\bigl\{\|\delta_{n+k}\|,\ \|\delta_{n+k-1}\|,\ldots,\|\delta_{n+1}\|\bigr\}\leq(1+\widetilde{C}h)\|\Delta_{n}\|_{\infty}.$ (27)
+Iterating and using $nh\leq T$ yields
+$\|\Delta_{n}\|_{\infty}\leq(1+\widetilde{C}h)^{n}\|\Delta_{0}\|_{\infty}\leq e^{\widetilde{C}T}\|\Delta_{0}\|_{\infty},\qquad 0\leq nh\leq T,$
+which completes the proof. ∎
+Moreover, the properties of EKahan scheme established for systems with a cubic Hamiltonian can be extended to higher-degree polynomial Hamiltonians, as summarized next.
+Theorem 4.
+For semilinear system (2) with polynomial potential $U(x)$ of degree $k+2$, the EKahan scheme (19) is symmetric.
+Proof.
+From equation (19) we get
+$x_{n+k}-e^{khA}x_{n}=kh\,\phi(khA)\,Q\,\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k}),\qquad A=QM.$ (28)
+Using $\phi(-Z)=e^{-Z}\phi(Z)$ and $e^{-khA}=(e^{khA})^{-1}$, we can rewrite (28) as
+$x_{n}-e^{-khA}x_{n+k}=-kh\,\phi(-khA)\,Q\,\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k}).$
+Since $\nabla_{\mathrm{K}}U$ is symmetric in its arguments, $\nabla_{\mathrm{K}}U$ will not change if we reverse the sequence $(x_{n},\ldots,x_{n+k})$. Therefore, the method is symmetric. ∎
+Lemma 4.
+Let $A\in\mathbb{R}^{d\times d}$ and $g\in C^{2}([0,kh];\mathbb{R}^{d})$. For all $t\in[0,kh]$, we claim
+$\int_{0}^{kh}e^{(kh-s)A}g(s)\mathrm{d}s=kh\,\phi(khA)\,g(kh/2)+R,\qquad\|R\|\leq Ch^{3},$
+where $C$ depends on $k$, $M_{E}$, $\|A\|$, and $\max\limits_{0\leq s\leq kh}\bigl(\|g(s)\|+\|g^{\prime}(s)\|+\|g^{\prime\prime}(s)\|\bigr)$.
+Lemma 5.
+Assume $U$ is homogeneous and let $x(t)$ be a smooth solution on $t_{n},t_{n}+kh]$.
+Set $\tilde{x}_{n+i}:=x(t_{n}+ih)$ and $t_{n+\frac{k}{2}}:=t_{n}+\frac{k}{2}h$.
+Then there exists $C>0$, independent of $h$, such that
+$\bigl\|\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})-\nabla U\bigl(x(t_{n+\frac{k}{2}})\bigr)\bigr\|\leq Ch^{2}.$
+Lemma 6.
+Let $x(t)$ solve (2) with $U$ homogeneous of degree $k+2$ and set $\tilde{x}_{n+i}=x(t_{n}+ih)$. Then the local defect of (19) satisfies
+$\|\tau_{n+k}\|\leq C_{\mathrm{loc}}\,h^{3},\qquad t_{n}\leq T,$
+for a constant $C_{\mathrm{loc}}$ independent of $h$.
+Proof.
+By the variation of constants formula for $\dot{x}=Ax+Q\nabla U(x)$, we have
+$\tilde{x}_{n+k}=e^{khA}\tilde{x}_{n}+\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s.$
+Hence
+$\tau_{n+k}=\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s-kh\,\phi(khA)\,Q\,\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k}).$
+Add and subtract the exponential midpoint quantity $kh\,\phi(khA)\,Q\,\nabla U(x(t_{n+\frac{k}{2}}))$:
+$\begin{split}\tau_{n+k}&=\underbrace{\Bigl[\int_{0}^{kh}e^{(kh-s)A}Q\nabla U(x(t_{n}+s))\mathrm{d}s-kh\,\phi(khA)\,Q\,\nabla U(x(t_{n+\frac{k}{2}}))\Bigr]}_{=:I_{1}}\\ &+\underbrace{kh\,\phi(khA)\,Q\Bigl[\nabla U(x(t_{n+\frac{k}{2}}))-\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})\Bigr]}_{=:I_{2}}.\end{split}$
+For $I_{1}$, apply Lemma 4 with $g(s)=Q\nabla U(x(t_{n}+s))$ to get $\|I_{1}\|\leq Ch^{3}$. For $I_{2}$, use the boundedness of $\phi(khA)$ on $[0,T]$ and Lemma 5 to obtain
+$\|I_{2}\|\leq kh\,\|\phi(khA)\|\,\|Q\|\,\bigl\|\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})-\nabla U(x(t_{n+\frac{k}{2}}))\bigr\|\leq C\,h\cdot h^{2}=Ch^{3}.$
+Combining the two bounds yields $\|\tau_{n+k}\|\leq C_{\mathrm{loc}}h^{3}$. ∎
+Lemma 7.
+Assume the exact solution stays in a bounded set $\mathcal{D}$ on $[0,T]$ and $U$ is polynomial.
+Then there exist $h_{0}>0$ and $C>0$ such that for all $0<h\leq h_{0}$ the $k$-step EKahan update defines a unique one-step map $Y_{n+1}=\Phi_{h}(Y_{n})$ on $\mathcal{D}^{k}$, and
+$\|\Phi_{h}(Y)-\Phi_{h}(Z)\|_{\infty}\leq(1+Ch)\,\|Y-Z\|_{\infty},\qquad\forall\,Y,Z\in\mathcal{D}^{k},$
+Theorem 5.
+Under the assumptions of Lemma 7, suppose the starting values $x_{0},\ldots,x_{k-1}$ satisfy $\|x_{j}-\tilde{x}_{j}\|\leq Ch^{2}$ for $j=0,\ldots,k-1$. Then for $t_{n}=nh\leq T$ and $h\leq h_{0}$,
+$\max_{0\leq n\leq N}\|x_{n}-\tilde{x}_{n}\|\leq C_{T}\,h^{2},$
+i.e., the $k$-step EKahan scheme is globally second order on finite time intervals.
+Proof.
+Inserting the exact solution values $\tilde{Y}_{n}$ into the scheme gives
+$\tilde{Y}_{n+1}=\Phi_{h}(\tilde{Y}_{n})+E_{n+1},\qquad\|E_{n+1}\|\leq C_{\mathrm{loc}}h^{3}.$
+where $E_{n}$ contains only the last component defect $\tau_{n+k}$ from Lemma 6.
+Subtracting from $Y_{n+1}=\Phi_{h}(Y_{n})$ and using Lemma 7 yields the following recursion for the extended error $\Delta_{n}=Y_{n}-\tilde{Y}_{n}$
+$\|\Delta_{n+1}\|_{\infty}\leq(1+Ch)\|\Delta_{n}\|_{\infty}+C_{\mathrm{loc}}h^{3}.$
+Applying the discrete Gronwall inequality gives, for $t_{n}\leq T$,
+$\|\Delta_{n}\|_{\infty}\leq e^{CT}\|\Delta_{0}\|_{\infty}+\frac{C_{\mathrm{loc}}}{C}\,(e^{CT}-1)\,h^{2}.$
+Since $\|\Delta_{0}\|_{\infty}=\max\limits_{0\leq j\leq k-1}\|x_{j}-\tilde{x}_{j}\|=O(h^{2})$ by assumption, the claim follows. ∎
 Corollary 2.
-For the semilinear system (2), any exponential integrator with the form
+For the semilinear system (2), any exponential integrator of the form
 $x_{n+k}=e^{khA}x_{n}+kh\phi(khA)\,Q\,\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k}),$
-where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretiztion of the true gradient $\nabla U(x_{n})$, satisfies the following equation
+where $\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretization of the true gradient $\nabla U(x_{n})$, satisfies
 $\frac{1}{2k}x_{n+k}^{\mathsf{T}}Mx_{n+k}-\frac{1}{2k}x_{n}^{\mathsf{T}}Mx_{n}+\frac{1}{k}(x_{n+k}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})=0.$
-The proof of this corollary follows similarly as the proof of Lemma 3.
 Corollary 3.
-For a homogeneous function $U(x)$, EKahan scheme (16) nearly preserves the exact energy and has the following error estimation:
-$H_{n+1}-H_{n}=$ $\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})$
-$+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})$
-$-2\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}),$
-where $H_{n}=\frac{1}{2k}\sum\limits_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$ is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\cdot,\dots,\cdot,\cdot)$ is a symmetric $k+2$ linear function defined via the $k+2$ tensor associated with the homogeneous function $U(x)$ and satisfies $\bar{U}(x,x,\dots,x,x)=U(x)$.
-The proof of this corollary proceeds analogously to that of Theorem 2.
-For polynomial nonlinear functions $U(x)$ of order greater than three, the energy error can no longer be expressed in the simple form presented in Theorem 2; nevertheless, Corollary 3 demonstrates that the energy error over a single time step can be effectively controlled by the corresponding change in the numerical solution, thereby ensuring the boundedness of the energy error.
+For a homogeneous function $U(x)$ of degree $k+2$, EKahan scheme (19) satisfies the stepwise identity
+$\begin{split}H_{n+1}-H_{n}=&\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})\\ &+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})\\ &-2\bar{U}\Bigl(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}\Bigr),\end{split}$ (29)
+where
+$H_{n}=\frac{1}{2k}\sum_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$
+is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\dots,\cdot)$ is a symmetric $(k+2)$-linear form associated with $U$ satisfying $\bar{U}(x,\dots,x)=U(x)$.
 Remark 2.
-For a nonhomogeneous polynomial $U(x)$, we can extend it to a homogeneous function by introducing an auxiliary variable $x^{0}$ to the original vector $x$.
-Denote the extended variable as $\bar{x}=(x^{0},x^{1},\cdots,x^{m})^{\mathsf{T}}$, where $x^{i}$, $i=1,\cdots,m$, are the components of the vector $x$.
-The nonhomogeneous cubic polynomial $U(x)$ is extended to the homogeneous function $\bar{U}(x^{0},x^{1},\cdots,x^{m})$ such that $\bar{U}(1,x^{1},\cdots,x^{m})=U(x)$.
-Furthermore, to preserve the structural properties of the original system in the extended formulation, we also augment the skew-symmetric matrix $Q$ to $\bar{Q}$ and the symmetric matrix $M$ to $\bar{M}$ by introducing an extra row and column of zero elements.
-Notably, $\bar{Q}$ remains skew-symmetric.
-Furthermore, applying the EKahan method to the modified differential equation
+For a nonhomogeneous polynomial $U(x)$, we homogenize it by introducing an auxiliary variable $x^{0}$ and an extended state $\bar{x}=(x^{0},x^{1},\ldots,x^{m})^{\mathsf{T}}$ so that $\bar{U}(1,x)=U(x)$.
+To preserve the structural properties of the original system, we also augment $Q$ to $\bar{Q}$ and $M$ to $\bar{M}$ by adding a zero row and column.
+Notably, $\bar{Q}$ remains skew-symmetric and $\bar{M}$ remains symmetric.
+Furthermore, applying the EKahan method to
 $\dot{\bar{x}}(t)=\bar{Q}\bar{M}\bar{x}(t)+\bar{Q}\nabla\bar{U}(\bar{x}(t)),\quad\bar{x}(t_{0})=[1,x_{0}^{\mathsf{T}}]^{\mathsf{T}},$
-produces the same update for $x$ as applying the EKahan method directly to the original reduced system (2).
-Therefore, all results previously established for systems with homogeneous polynomial nonlinearities extend naturally to those involving nonhomogeneous polynomial terms.
+produces the same update for $x$ as applying the EKahan method directly to the original system (2).
+Therefore, all results previously established for systems with homogeneous polynomial Hamiltonians extend naturally to those involving nonhomogeneous polynomials.
 3 Numerical test
@@
   • LIEEP: the linearly implicit energy-preserving exponential scheme illustrated in [22].
-We define the global error by
-$\mathcal{E}_{G}=\max\limits_{n\geq 0}\|y_{n}-y(t_{n})\|,$
-and the energy error at time $t_{n}=t_{0}+nh$ by
-$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert,$
-where $y(t_{n})$ is the reference exact solution and $H_{n}$ is the discrete energy defined in Theorem 2.
+We define the global error at time $t_{n}=t_{0}+nh$ as
+$\mathcal{E}_{G}=\max_{n\geq 0}\|y_{n}-y(t_{n})\|,$
+where $y(t_{n})$ is the reference exact solution. For EKahan, we measure the residual of the stepwise energy-increment identity by
+$e_{n}=H_{n+1}-H_{n}-G_{n},$
+where $G_{n}=U(x_{n+1}-x_{n})$ for cubic Hamiltonians and $H_{n}$ is the discrete energy defined in Theorem 2.
+For higher-degree polynomial Hamiltonians, $G_{n}$ is given by the right-hand side of equation (29) and $H_{n}$ is defined as in Corollary 3.
+The energy error is defined as
+$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert$. (30)
 The reference exact solution is obtained using the 6-order continuous Runge-Kutta (CRK) method [14] with the form
@@
 These starting values are obtained using the aforementioned reference exact solution.
+The matrix exponential functions $e^{khQM}$ and $\phi(khQM)$ are computed using the MATLAB package presented in [2] when no explicit form is given, where Padé approximations are employed.
+We solve the linear systems arising from each step of the linearly implicit schemes (EKahan, Kahan and LIEEP) by MATLAB’s backslash operator.
+The nonlinear updates required for the reference solutions and the EAVF numerical solutions are computed via a fixed-point iteration, terminated when the 2-norm of the difference between successive iterates falls below a tolerance of $10^{-14}$.
+We also tested tolerances from $10^{-14}$ to $10^{-6}$ for EAVF and observed that EKahan is consistently the most efficient method across our test problems, except for the FPU system with $p=2$, where EKahan is not as efficient as EAVF when a bigger tolerance is considered.
 All numerical experiments were conducted using MATLAB (R2022b) on a 2020 MacBook Air equipped with an Apple M1 processor and 16 GB of unified memory.
-Henon-Heiles system
-The Henon-Heiles model [32] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
-$\left\{\begin{array}[]{ll}{\ddot{q}_{1}=-q_{1}-2Dq_{1}q_{2}},\\ {\ddot{q}_{2}=-q_{2}-Dq_{1}^{2}+Cq_{2}^{2}},\\ \end{array}\right.$
+Hénon-Heiles system
+The Hénon-Heiles model [11] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
+[…]
 where the independent variable $t\in[0,T]$.
@@
 By introducing the momentum variables $\dot{q}_{1}=p_{1},\;\dot{q}_{2}=p_{2}$, the system can be rewritten in the Hamiltonian form
-$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$ (17)
+$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$
 where $y=(q_{1},q_{2},p_{1},p_{2})^{\mathsf{T}}$, $Q$ is the canonical skew-symmetric matrix, $M=I_{4}$ denotes the $4\times 4$ identity matrix, and the potential function is defined as $U(y)=q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}$.
-Applying the EKahan scheme to equation (17) gives
-$y_{n+1}=e^{hQM}y_{n}-h\phi(hQM)Q[\frac{1}{2}\nabla{U}(y_{n})-2\nabla U(\frac{y_{n+1}+y_{n}}{2})+\frac{1}{2}\nabla U(y_{n+1})],$
-and we can get the accurate form of the exponential function
+The exponential function can be computed explicitly by
 $e^{hQM}=\begin{pmatrix}\cos(h)I_{2}&\sin(h)I_{2}\\ -\sin(h)I_{2}&\cos(h)I_{2}\end{pmatrix}$.
@@
 Figure 1(a) demonstrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded, varying from $10^{-18}$ to $10^{-7}$.
-Figure 1(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
-Figure 1: Energy behaviour for the Henon-Heiles equation with $T=100$ and $h_{0}=0.02$.
-Left: energy error computed by all schemes; right: the error of the true and deduced energy deviation by EKahan method shown in Theorem 2.
+Figure 1(b) verifies the result about the energy deviation at each time step by EKahan shown in Theorem 2.
+Furthermore, we verified numerically that the per-step energy increment of EKahan scales as $O(h^{3})$, consistent with the stepwise identity in Theorem 2, and we omitted the corresponding figure due to space limitations.
+Figure 1: Hénon-Heiles equation with $T=100$ and $h_{0}=0.02$. Left: energy error defined in (30); right: the residual of equation (16) in Theorem 2.
 In Figure 2, we consider the global errors at $T=100$ and the computational cost using different step sizes.
@@
 (b) efficiency plot
-Figure 2: Global error and computational time of different methods for the Henon-Heiles equation with $T=100$ and different time step sizes.
+Figure 2: Global error and computational time of different methods for the Hénon-Heiles equation with $T=100$ and different time step sizes.
 FPU system with a cubic Hamiltonian
@@
 where the independent variables $(x,t)\in[0,L]\times[0,T]$, the parameter $\epsilon>0$, the internal damping coefficient $\beta\geq 0$, and the external damping coefficient $\gamma\geq 0$.
-By introducing $\frac{\partial u}{\partial t}=v$, we can reformulate the second-order partial differential equation into the following first-order differential equations
-$\left\{\begin{array}[]{ll}{\partial_{t}u=v,}\\ {\partial_{t}v=\beta\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}u}{\partial x^{2}}(1+\epsilon(\frac{\partial u}{\partial x})^{p})-\gamma v-m^{2}u}.\\ \end{array}\right.$
-Setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
+By introducing $\frac{\partial u}{\partial t}=v$ and setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
 $\frac{\partial y}{\partial t}=\mathcal{Q}\frac{\delta\mathcal{H}}{\delta y},$
@@
 where
-$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u_{j+1}-u_{j}}{\Delta x})^{3}.$
+$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u^{j+1}-u^{j}}{\Delta x})^{3}.$
 We set $m=0$, $\epsilon=\frac{3}{4}$, $L=128$, $T=100$, the spatial discretization to be $\Delta x=1$ and the time step size to be $h_{i}=1/2^{i}$, $i=1,2,\ldots,4$.
 The initial conditions are given by
-$u_{j}(0)=q_{j}(0),\quad v_{j}(0)=\dot{q}_{j}(0),$
-where the function $q_{j}(t)$ takes the form
-$q_{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
+$u^{j}(0)=q^{j}(0),\quad v^{j}(0)=\dot{q}^{j}(0),$
+where the function $q^{j}(t)$ takes the form
+$q^{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
 with $\alpha=0.1$.
-We consider three different settings of damping coefficients: i) a conservative system with $\gamma=0$ and $\beta=0$, ii) a dissipative system with $\gamma=0.1$ and $\beta=0$, and iii) a dissipative system with $\gamma=0$ and $\beta=2$.
-Figure 3 shows that EKahan method preserves the energy structure of the equation in both conservative and dissipative settings:
-  1. i)
+We consider three different settings of damping coefficients: 1) a conservative system with $\gamma=0$ and $\beta=0$, 2) a dissipative system with $\gamma=0.1$ and $\beta=0$, and 3) a dissipative system with $\gamma=0$ and $\beta=2$.
+We emphasize that our theoretical results are derived for the conservative Hamiltonian case, where the stepwise energy-increment identity in Theorem 2 applies.
+The two dissipative settings are included as additional numerical tests to assess whether EKahan captures the expected energy-decay behavior when damping is present.
+Figure 3 summarizes the observed energy behavior:
   • the energy error of all methods stays bounded and oscillated in the undamped case in Figure 3(a),
-  2. ii)
   • the per-step energy deviation under EKahan discretization shown in Theorem 2 is validated in Figure 3(b) ,
-  3. iii)
-    the energy of all methods drop along time for systems with a positive internal or external damping coefficient in Figure 3(c) and 3(d).
+  • in the damped cases ($\gamma>0$ or $\beta>0$), the computed energy decreases over time, consistent with the dissipative nature of the underlying dynamics; see Figures 3(c) and 3(d).
 (a) $\gamma=0$, $\beta=0$
@@
 (d) $\gamma=0$, $\beta=2$
-Figure 3: Energy behaviour for the FPU system with $p=1$, $T=100$, and $h_{2}=0.25$.
-Figure 3(a): the energy error computed by all schemes for the conservative system; Figure 3(b): the error of the true and deduced energy deviation by EKahan method shown in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy computed by all schemes for the two dissipative systems.
+Figure 3: FPU system with $p=1$, $T=100$, and $h_{2}=0.25$.
+Figure 3(a): the energy error defined in (30) computed by all schemes for the conservative system; Figure 3(b): the residual of equation (16) in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy defined in Theorem 2 of all schemes for the two dissipative systems.
 We test the convergence order and computational efficiency for all methods with those three settings of damping coefficients and the results are consistent.
 We report the results for the conservative equations in Figure 4.
-Figure 4(a) shows that Ekahan method is second-order and and Figure 4(b) shows that EKahan method is more efficient than all other methods.
-In addition, we observe that Kahan method is less accurate than any exponential integrator although it is shown to be more computationally effective than EAVF method.
+Figure 4(a) shows that EKahan is second-order and and Figure 4(b) shows that EKahan is more efficient than all other methods.
 In Figure 5, we illustrate the numerical solutions given by EKahan method.
-Figure 5(a) shows that EKahan method captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time.
-Figure 5(b) shows that EKahan method captures the decay of the wave profiles as energy dissipates internally.
+Figure 5(a) shows that EKahan captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time.
+Figure 5(b) shows that EKahan captures the decay of the wave profiles as energy dissipates internally.
 Figure 3(c) and 5(b) demonstrate that although EKahan is designed for conservative equations, it can also capture the dissipative nature for the dissipative systems.
@@
 (b) $\gamma=0,\beta=2$
-Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step need confirmation $h_{2}=0.25$.
+Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step $h_{2}=0.25$.
 FPU system with a fourth-order Hamiltonian
-We consider $p=2$, i.e., the potential $U(y)$ is a quartic homogeneous polynomial function.
+When $p=2$, the potential function is a quartic homogeneous polynomial with the form
+$U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{12}\left(\frac{u^{j+1}-u^{j}}{\Delta x}\right)^{4}.$
+Define the spatial discrete difference by
+$w^{j}=\frac{u^{j+1}-u^{j}}{\Delta x},\quad j=0,\ldots,N-1.$
+Following equation (18), the polarized gradient is given by
+$\left(\nabla_{\mathrm{K}}U(y_{n},y_{n+1},y_{n+2})\right)^{j}=\frac{\epsilon}{3\Delta x}\left(w^{j-1}_{n}w^{j-1}_{n+1}w^{j-1}_{n+2}-w^{j}_{n}w^{j}_{n+1}w^{j}_{n+2}\right)$
+for $j=0,\ldots,N-1$. The EKahan scheme then takes the form
+$y_{n+2}=e^{2hQM}y_{n}+2h\phi(2hQM)Q\nabla_{\mathrm{K}}U(y_{n},y_{n+1},y_{n+2}).$
 We employ identical initial boundary condition and computational parameters to the $p=1$ case, with the exception of setting $\epsilon=100$.
@@
 Figure 6(b) verifies the formula of the energy deviation by EKahan method shown in Corollary 3.
-Figure 6: Energy behaviour for the FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$.
-Left: the energy error computed by all schemes; right: the error of the true and deduced energy deviation computed by EKahan method shown in Corollary 3.
+Figure 6: FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$. Left: the energy error (30) of all schemes; right: the residual of equation (29) in Corollary 3 by EKahan.
 In Figure 7, we consider the global error and the computational cost using different time step sizes.
@@
 For $p=1$, the equation admits the following Hamiltonian structure:
-$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$
+$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$ (31)
 To discretize the problem, we use the following discrete energy
-$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u_{i,j})^{2}+(\delta_{x}^{-}u_{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u_{i,j})^{2}+(\delta_{y}^{-}u_{i,j})^{2}}{4}+\frac{u_{i,j}^{3}}{6}\right)\Delta x\Delta y.$
+$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u^{i,j})^{2}+(\delta_{x}^{-}u^{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u^{i,j})^{2}+(\delta_{y}^{-}u^{i,j})^{2}}{4}+\frac{(u^{i,j})^{3}}{6}\right)\Delta x\Delta y.$
 Here, $\delta_{x}^{+}$ and $\delta_{x}^{-}$, similarly $\delta_{y}^{+}$ and $\delta_{y}^{-}$, denote the forward and backward finite difference operators in the $x$- and $y$-directions, respectively.
-First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives, respectively.
-The corresponding discrete differential operators are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$.
+First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives in (31), respectively, and the corresponding matrices are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$.
 Using these notations, the semi-discrete system can be compactly written as
-$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U.^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
+$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
 where the vector $U\in\mathbb{R}^{(N_{x}-1)\times(N_{y}-1)}$ is flattened from $u_{i,j}$ by stacking columns successively, giving priority to the $y$-direction, and $U^{2}$ denotes the element-wise square of $U$.
 Here, $N_{x}$ and $N_{y}$ are the grid numbers in each spatial direction.
-Numerical simulations are carried out on the spatial domain $[0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions.
+Numerical simulations are carried out on the spatial domain $0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions.
 The spatial discretization employs a uniform grid with $N=33$ points in each spatial direction, resulting in a mesh size of $\Delta x=\Delta y=L/(N-1)$.
-For time integration, different time steps $h_{i}=0.01/2^{i+1},\;i=1,2,3,4$ are adopted.
-We consider the initial condition
+We consider different time steps $h_{i}=0.01/2^{i+1},\;i=1,\cdots,4$, and the initial condition
 $u(0,x,y)=\sqrt{2}[\sin(\frac{2\pi x}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi x}{L}+\frac{\pi}{4})]\times[\cos(\frac{2\pi y}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi y}{L}+\frac{\pi}{3})].$
 Figure 8(a) demonstrates that the energy errors of all linearly implicit methods remain bounded and oscillated.
-Figure 8(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
-Figure 8: Energy behaviour for the Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$.
-Left: the energy error computed by EKahan, Kahan and LIEEP schemes; right: the error of the true and deduced energy deviation by EKahan shown in Theorem 2.
-In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$.
-Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to Kahan method and LIEEP method.
+Figure 8(b) confirms the stepwise energy-increment identity for EKahan established in Theorem 2.
+Figure 8: Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$. Left: the energy error defined in (30); right: the residual of equation (16) in Theorem 2.
 (a) order plot
@@
 Figure 9: Global error and computational time of different methods for the Zakharov-Kuznetsov equation with $T=8$ and different time step sizes $h_{i}$.
+Figure 10: Evolution to bell-shaped solitons in Zakharov-Kuznetsov equation, resolved with time step $h_{1}=0.0025$. Left: initial condition; right: soliton structure at $t=8$.
+In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$.
+Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to the other methods.
 Figure 10 illustrates the solution of the Zakharov-Kuznetsov equation.
@@
 Despite the significant structural changes in the solution profile, the total energy of the system remains nearly constant throughout the simulation, demonstrating the excellent long-time stability.
-Figure 10: Evolution to bell-shaped solitons in Zakharov-Kuznetsov equation, resolved with time step $h_{1}=0.0025$. Left: initial condition; right: soliton structure at $t=8$.
 4 Conclusion
-In this work, we proposed a linearly implicit, structure-preserving exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities by combining Kahan’s method with exponential integration techniques.
-The resulting scheme, EKahan method, successfully balances computational efficiency with the preservation of critical geometric structures, such as symmetry and near-conservation of energy.
-Compared to fully implicit exponential integrators, such as the EAVF method, EKahan requires only the solution of a single linear system per time step, significantly reducing computational costs.
-Although not all geometric properties of Kahan’s method are exactly inherited by EKahan method, we rigorously proved that EKahan maintains symmetry and a bounded and oscillatory energy error and achieves second-order convergence.
-Through extensive numerical experiments on benchmark systems, we verified the method’s stability, long-term accuracy, and computational superiority.
+In this work, we proposed a linearly implicit structure-preserving exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities by combining the Kahan-type discretization with exponential integration techniques.
+Although not all geometric properties of Kahan’s method are exactly inherited by EKahan, the resulting scheme balances computational efficiency with key geometric features, including symmetry and exact stepwise energy-increment identities.
+Together with our second-order convergence result, these identities provide a computable near conservation of energy characterization in terms of numerical increments, rather than implying exact energy conservation or symplecticity for general nonlinear systems.
+Compared to the fully implicit energy-preserving exponential integrator EAVF method, EKahan requires only the solution of a single linear system per time step, significantly reducing computational costs.
+Through extensive numerical experiments on benchmark systems, we verified the method’s long-term accuracy and computational competitiveness.
 In particular, EKahan was shown to outperform not only the fully implicit scheme but also other linearly implicit methods, LIEEP and Kahan method, highlighting its enhanced efficiency among structure-preserving approaches.
-Furthermore, we provided the generalization of EKahan to problems involving high-order polynomial Hamiltonians, thereby broadening its applicability to a wider class of physical models.
+We also tested the method in representative damped regimes to assess its robustness beyond the conservative setting, where EKahan captures the expected energy-decay trend observed in the dynamics.
+Furthermore, we provided the generalization of EKahan to problems involving higher-degree polynomial Hamiltonians, thereby broadening its applicability to a wider class of physical models.
 Overall, the EKahan method offers a compelling tool for the long-time integration of semilinear Hamiltonian systems.
 Future work may explore extending this framework to adaptive step-size strategies, and constructing higher-order efficient structure-preserving exponential integrators.
+Appendix
+Lemma 2.
+For the semilinear system (2), any exponential integrator of the form
+$x_{n+1}=e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1}),$
+satisfies the following equation
+$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n},x_{n+1})=0,$
+where $\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
+Proof.
+We first assume that the matrix $M$ is not singular.
+Denoting by $V=hA$ and $\breve{\nabla}U=M^{-1}\hat{\nabla}U$ and using $\phi(V)V=e^{V}-I$, we get the following equation after replacing $x_{n+1}$ by $e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1})$
+$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}{x_{n}}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1})$
+$=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+\frac{1}{2}h^{2}\hat{\nabla}U^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U$
+$+h{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}-I)\hat{\nabla}U+h{\hat{\nabla}U}^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}\hat{\nabla}U$
+$=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}(e^{V}-I)^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U$
+$+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{\breve{\nabla}U}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U$
+$=\frac{1}{2}(x_{n}+\breve{\nabla}U)^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)(x_{n}+\breve{\nabla}U)+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-Me^{V})\breve{\nabla}U$
+$=0.$
+The last equation follows from the fact that ${e^{V}}^{\mathsf{T}}Me^{V}-M$ and ${e^{V}}^{\mathsf{T}}M-Me^{V}$ are skew symmetric.
+When $M$ is singular, we follow the idea in [23], introducing a series of symmetric and nonsingular matrices $M_{\delta}$ which converges to $M$ when $\delta$ approaches zero to complete the proof.
+∎
+Corollary 1.
+Let the potential $U$ be a general cubic polynomial with homogeneous decomposition as
+$U(x)=U_{3}(x)+U_{2}(x)+U_{1}(x)+U_{0},$
+where $U_{j}$ is homogeneous of degree $j$. Then the EKahan scheme (3) satisfies
+$H_{n+1}-H_{n}=U_{3}(x_{n+1}-x_{n}),$
+showing that the stepwise energy defect depends only on the cubic component.
+Proof.
+By Lemma 2 with $\hat{\nabla}U=\nabla_{\mathrm{K}}U$ and linearity of $\nabla_{\mathrm{K}}$ in $U$ and the linearity of $\nabla_{\mathrm{K}}$ regarding $U$, it suffices to treat each homogeneous part.
+For the homogeneous cubic part $U_{3}$, Theorem 2 yields
+$U_{3}(x_{n+1})-U_{3}(x_{n})-(x_{n+1}-x_{n})^{\top}\nabla_{\mathrm{K}}U_{3}(x_{n},x_{n+1})=U_{3}(x_{n+1}-x_{n}).$
+For $U_{2}$ and $U_{1}$, the Kahan polarization coincides with the exact discrete gradient identity
+$(x_{n+1}-x_{n})^{\top}\nabla_{\mathrm{K}}U_{j}(x_{n},x_{n+1})=U_{j}(x_{n+1})-U_{j}(x_{n}),\qquad j=1,2,$
+and $U_{0}$ contributes nothing.Therefore, only the cubic component contributes to the energy drift. ∎
+Lemma 3.
+Let $U$ be a polynomial of degree $k+2$.
+Then for each $n$ there exist a matrix $B_{n}\in\mathbb{R}^{d\times d}$ and a vector $\ell_{n}\in\mathbb{R}^{d}$, depending only on $(x_{n},\dots,x_{n+k-1})$, such that
+$\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k})=B_{n}x_{n+k}+\ell_{n}.$ (32)
+Consequently, (19) is equivalent to the linear system
+$\Bigl(I-kh\,\phi(khQM)\,QB_{n}\Bigr)x_{n+k}=e^{khQM}x_{n}+kh\,\phi(khQM)\,Q\ell_{n}.$ (33)
+In particular, if $U$ is homogeneous, then $\ell_{n}=0$.
+Proof.
+Fix $(x_{n},\dots,x_{n+k-1})$ and regard $\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k})$ in (18) as a function of the last argument $x_{n+k}$.
+Since $U$ is a polynomial, $\nabla U$ is a polynomial vector field, and the polarization formula (18) is obtained from the symmetric polarization of each homogeneous component of $\nabla U$.
+Therefore $\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k})$ is linear in $x_{n+k}$, and we may write
+$\ell_{n}:=\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k-1},0),\qquad B_{n}x:=\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k-1},x)-\ell_{n},$
+which yields (32).
+Substituting (32) into (19) gives (33).
+Finally, if $U$ is homogeneous of degree $k+2$, then $\nabla U$ is homogeneous of degree $k+1$ and its polarization is (k+1)(k\!+\!1)-linear, so $\nabla_{\mathrm{K}}U(x_{n},\dots,x_{n+k-1},0)=0$, i.e., $\ell_{n}=0$.
+∎
+Lemma 4.
+Let $A\in\mathbb{R}^{d\times d}$ and $g\in C^{2}([0,kh];\mathbb{R}^{d})$. For all $t\in[0,kh]$, we claim
+$\int_{0}^{kh}e^{(kh-s)A}g(s)\,ds=kh\,\phi(khA)\,g(kh/2)+R,\qquad\|R\|\leq Ch^{3},$
+where $C$ depends on $k$, $M_{E}$, $\|A\|$, and $\max\limits_{0\leq s\leq kh}\bigl(\|g(s)\|+\|g^{\prime}(s)\|+\|g^{\prime\prime}(s)\|\bigr)$.
+Proof.
+Set $s_{*}=kh/2$ and $F(s)=e^{(kh-s)A}g(s)$. Then $F\in C^{2}$ and by the classical midpoint rule with Peano remainder, we have
+$\int_{0}^{kh}F(s)\,ds=kh\,F(s_{*})+R_{1},\qquad\|R_{1}\|\leq\frac{(kh)^{3}}{24}\max_{0\leq s\leq kh}\|F^{\prime\prime}(s)\|.$ (34)
+A direct differentiation gives
+$F^{\prime\prime}(s)=e^{(kh-s)A}\Bigl(A^{2}g(s)-2Ag^{\prime}(s)+g^{\prime\prime}(s)\Bigr),$
+hence
+$\max_{0\leq s\leq kh}\|F^{\prime\prime}(s)\|\leq M_{E}\Bigl(\|A\|^{2}\max\|g\|+2\|A\|\max\|g^{\prime}\|+\max\|g^{\prime\prime}\|\Bigr).$
+Therefore, there exist $C_{1}$ such that $\|R_{1}\|\leq C_{1}h^{3}$.
+Defining a function $D(z)=e^{z/2}-\phi(z)$, we have $D(z)=-\frac{z^{2}}{24}+O(z^{3})$ as $z\to 0$. Thus, there exists $C_{2}$, depending on $k$, $M_{E}$, $\|A\|$, such that
+$\|D(khA)\|\leq C_{2}h^{2};$
+and therefore, there exists $C_{3}$ such that
+$kh\,e^{(kh/2)A}g(s_{*})=kh\,\phi(khA)g(s_{*})+kh\,D(khA)g(s_{*})=kh\,\phi(khA)g(kh/2)+R_{2},$
+and $\|R_{2}\|\leq C_{3}h^{3}$. Combining with (34) yields the claim with $R=R_{1}+R_{2}$, and $C=C_{1}+C_{3}$. ∎
+Lemma 5.
+Assume $U$ is homogeneous of degree $k+2$ and let $x(t)$ be a smooth solution on $t_{n},t_{n}+kh]$.
+Set $\tilde{x}_{n+i}:=x(t_{n}+ih)$ and $t_{n+\frac{k}{2}}:=t_{n}+\frac{k}{2}h$.
+Then there exists $C>0$ (depending on $U$ and bounds of $x(t)$ on the macro-interval) such that
+$\bigl\|\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})-\nabla U\bigl(x(t_{n+\frac{k}{2}})\bigr)\bigr\|\leq Ch^{2}.$
+Proof.
+Let $\bar{x}:=x(t_{n+\frac{k}{2}})$ and write $\tilde{x}_{n+i}=\bar{x}+\delta_{i}$. A Taylor expansion of $x(t)$ around $t_{n+\frac{k}{2}}$ gives
+$\delta_{i}=(i-\tfrac{k}{2})h\,\dot{x}(t_{n+\frac{k}{2}})+O(h^{2}),\qquad i=0,\ldots,k,$
+hence $\|\delta_{i}\|=O(h)$ and moreover
+$\sum_{i=0}^{k}\delta_{i}=O(h^{2}),\quad\text{since}\quad\sum_{i=0}^{k}(i-\tfrac{k}{2})=0.$
+For homogeneous $U$, the polarized gradient is symmetric and multilinear in its $(k+1)$ arguments; in particular, $\nabla_{\mathrm{K}}U(\bar{x},\ldots,\bar{x})=\nabla U(\bar{x})$.
+Using multilinearity around the diagonal point $(\bar{x},\ldots,\bar{x})$, we expand
+$\nabla_{\mathrm{K}}U(\bar{x}+\delta_{0},\ldots,\bar{x}+\delta_{k})=\nabla_{\mathrm{K}}U(\bar{x},\ldots,\bar{x})+\sum_{i=0}^{k}\mathcal{L}(\delta_{i})+O(h^{2}),$
+where $\mathcal{L}(\cdot)$ denotes the same linear map (by symmetry) obtained by replacing one argument $\bar{x}$ with the increment. Consequently,
+$\Bigl\|\sum_{i=0}^{k}\mathcal{L}(\delta_{i})\Bigr\|\leq C_{1}\Bigl\|\sum_{i=0}^{k}\delta_{i}\Bigr\|=O(h^{2}),$
+and the remaining terms contain at least two increments and are therefore also $O(h^{2})$. Combining these estimates yields
+$\nabla_{\mathrm{K}}U(\tilde{x}_{n},\ldots,\tilde{x}_{n+k})=\nabla U(\bar{x})+O(h^{2}),$
+which concludes the proof. ∎
+Lemma 6.
+Assume the exact solution stays in a bounded set $\mathcal{D}$ on $[0,T]$ and $U$ is polynomial.
+Then there exist $h_{0}>0$ and $C>0$ such that for all $0<h\leq h_{0}$ the $k$-step EKahan update defines a unique one-step map $Y_{n+1}=\Phi_{h}(Y_{n})$ on $\mathcal{D}^{k}$, and
+$\|\Phi_{h}(Y)-\Phi_{h}(Z)\|_{\infty}\leq(1+Ch)\,\|Y-Z\|_{\infty},\qquad\forall\,Y,Z\in\mathcal{D}^{k},$
+where $\|Y\|_{\infty}:=\max\limits_{0\leq j\leq k-1}\|y_{j}\|$ for $Y=(y_{0},\ldots,y_{k-1})\in(\mathbb{R}^{d})^{k}$.
+Proof.
+From Theorem 3, the update from $Y_{n}$ to $Y_{n+1}$ can be uniquely defined via solving (21). For given $Y$ and $Z$ in $\mathcal{D}^{k}$, we claim , we claim
+$\|\Phi_{h}(Y)-\Phi_{h}(Z)\|_{\infty}=\max\Bigl\{\max_{1\leq j\leq k-1}\|y_{j}-z_{j}\|,\ \|y_{k}-z_{k}\|\Bigr\}\leq(1+\widetilde{C}h)\,\|Y-Z\|_{\infty},$
+based on the estimation in (27). ∎
+Corollary 2.
+For the semilinear system (2), any exponential integrator of the form
+$x_{n+k}=e^{khA}x_{n}+kh\phi(khA)\,Q\,\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k}),$
+where $\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretization of the true gradient $\nabla U(x_{n})$, satisfies
+$\frac{1}{2k}x_{n+k}^{\mathsf{T}}Mx_{n+k}-\frac{1}{2k}x_{n}^{\mathsf{T}}Mx_{n}+\frac{1}{k}(x_{n+k}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})=0.$
+Proof.
+Assume first that $M$ is nonsingular. Set $V=khA$ and define $\breve{\nabla}U=M^{-1}\hat{\nabla}U$. Using $\phi(V)V=(e^{V}-I)$ and $A=QM$, we obtain
+$kh\,\phi(V)\,Q\,\hat{\nabla}U=(e^{V}-I)\breve{\nabla}U,\qquad x_{n+k}-x_{n}=(e^{V}-I)(x_{n}+\breve{\nabla}U).$
+A direct expansion gives
+$\frac{1}{2k}x_{n+k}^{T}Mx_{n+k}-\frac{1}{2k}x_{n}^{T}Mx_{n}+\frac{1}{k}(x_{n+k}-x_{n})^{T}\hat{\nabla}U$
+$=\frac{1}{2k}(x_{n}+\breve{\nabla}U)^{T}\bigl(e^{V^{T}}Me^{V}-M\bigr)(x_{n}+\breve{\nabla}U)+\frac{1}{2k}\breve{\nabla}U^{T}\bigl(e^{V^{T}}M-Me^{V}\bigr)\breve{\nabla}U,$
+which vanishes since $V^{\mathsf{T}}M+MV=0$ implies $e^{V^{\mathsf{T}}}Me^{V}=M$ and $e^{V^{T}}M-Me^{V}$ is skew-symmetric. The singular case follows the same as in Theorem 2. ∎
+Corollary 3.
+For a homogeneous function $U(x)$ of degree $k+2$, EKahan scheme (19) satisfies the stepwise identity
+$H_{n+1}-H_{n}=$ $\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})$
+$+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})$
+$-2\bar{U}\Bigl(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}\Bigr),$
+where
+$H_{n}=\frac{1}{2k}\sum_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$
+is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\dots,\cdot)$ is a symmetric $(k+2)$-linear form associated with $U$ satisfying $\bar{U}(x,\dots,x)=U(x)$.
+Proof.
+Applying Corollary 2 with $\hat{\nabla}U=\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k})$ gives
+$\frac{1}{2k}x_{n+k}^{T}Mx_{n+k}-\frac{1}{2k}x_{n}^{T}Mx_{n}=-\frac{1}{k}(x_{n+k}-x_{n})^{T}\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k}).$ (35)
+Recall that $U$ is homogeneous of degree $k+2$ and can be represented by a symmetric $(k+2)$-linear form $\bar{U}$ such that $U(x)=\bar{U}(x,\ldots,x)$.
+Moreover, the homogeneity implies the directional derivative identity
+$u^{T}\nabla U(z)=(k+2)\,\bar{U}(u,z,\ldots,z)\qquad\text{for all }u,z\in\mathbb{R}^{d}.$
+Since $\nabla_{\mathrm{K}}U$ is the $(k+1)$-linear polarization of $\nabla U$ and polarization commutes with linear functionals, it follows that for all $u\in\mathbb{R}^{d}$,
+$u^{T}\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k})=(k+2)\,\bar{U}(u,x_{n},\ldots,x_{n+k}).$
+Taking $u:=x_{n+k}-x_{n}$ and using multilinearity of $\bar{U}$ yields
+$\frac{1}{k}u^{T}\nabla_{\mathrm{K}}U(x_{n},\ldots,x_{n+k})=2\bar{U}\!\Bigl(x_{n},x_{n+1},\ldots,x_{n+k},\frac{u}{k}\Bigr)+\bar{U}(x_{n},x_{n+1},\ldots,x_{n+k},u).$ (36)
+By the definition of the discrete energy $H_{n}$ in the statement,
+$H_{n+1}-H_{n}$ $=\frac{1}{2k}(x_{n+k}^{T}Mx_{n+k}-x_{n}^{T}Mx_{n})+\bar{U}(x_{n+1},\ldots,x_{n+k},x_{n+1},x_{n+k})$
+$-\bar{U}(x_{n},\ldots,x_{n+k-1},x_{n},x_{n+k-1}).$
+Combining (35) and (36) and using the multilinearity and symmetry of $\bar{U}$ yields the stated increment formula. ∎